### v1

# Methods Paper Version 19

Abstract
We apply the Hulthén–Kohn method suggested by V. D. Efros [Phys. Rev. C 99, 034620 (2019)] for calculating various observables in the continuum and discrete spectrum using two-body interactions in single- and coupled-channel systems. This method is promising for many-body applications and ab initio description of nuclear reactions. We explore the convergence of phase shifts and wave functions as well as the location of $S$-matrix poles which enables obtaining both resonance and bound state parameters. We find that adopting wave functions from approximate bound-state solutions for the short-range components of basis wave functions leads to good convergence.
I Introduction
Over the last 25 years, various ab initio (from first principles) methods have been developed and extensively applied with the aim of describing the structure of atomic nuclei, including the Green’s function Monte Carlo method GFMC0 ; GFMC1 ; GFMC2 ; GFMC3 ; GFMC4 ; GFMC5 , the coupled cluster method CoesterOne ; CoesterTwo ; Kummel ; Wloh ; Hagen ; Hagen2022 ; ZHSun2023 , and the no-core shell model (NCSM) NCSMPRL ; NCSMProg . In principle, one can choose any basis within the Hilbert space to construct the eigenfunctions. For practical reasons, the harmonic oscillator basis is conventionally used in nuclear structure theory as it possesses several convenient features for applications to realistic many-body problems. The utility is such that the many-body problem is reduced within the M-scheme (in which all basis states have a fixed value of total angular momentum projection but not a fixed total angular momentum) to a sparse matrix eigenvalue problem that is solvable with leadership class supercomputers MFDnMaris2010 ; Euro-Par ; Aktulga ; NCSM-NTSE2014 ; Shao .
While there are many advantages in applications of those methods to nuclear structure problems, there are major technical difficulties in applying them to describe scattering states and reactions. While bound states can be calculated by ab initio approaches with a reasonable degree of accuracy for problems up to a nucleon number of approximately $A=20$ and for heavier nuclei around closed shells, successes in describing ab initio scattering in systems with $A>4$ are more limited CWJohnson . This is a pressing issue, since developing experimental techniques will benefit from accurate calculations of observables in nuclear reactions such as scattering cross sections, resonance energies $E_{r}$ and widths $\Gamma$, the boundaries of stability as either the proton or neutron number is increased VSoma , and properties of exotic systems like the tetraneutron and trineutron AMShirokov2016 ; 4nAIP ; JLi2019 ; 3narXiv .
In light of these needs, several ab initio methods have been applied to elastic scattering problems. For example, the Green’s function Monte Carlo has been used to obtain features of the $J^{\pi}=\frac{3}{2}^{-}$ and $\frac{1}{2}^{-}$ low-lying resonant states in $n\alpha$ scattering Nollett . Properties of nuclear resonances and reactions with light nuclei can also be studied within the no-core Gamow shell model Papadimitriou . In particular, this approach was successfully applied recently to the study of the tetraneutron and trineutron JLi2019 . Another successful approach is the resonating group method (RGM) in combination with the NCSM QuagOne ; QuagTwo ; Navratil2010 , which has been used to study nuclear scattering in systems as large as $A=17$ Navratil2010 , and its more extended version, the NCSM with continuum BaroniTwo ; BaroniOne ; Navr-scr ; NCSMCHebborn2022 ; NCSMCGysbers2023 .
An alternative approach to the studies of the continuum spectrum is the harmonic oscillator representation of scattering equations (HORSE) BangHORSE ; Zaitsev1998 . The HORSE method allows one to calculate the continuum wave function and the scattering phase shifts $\delta_{l}(E)$. It has already been successfully applied to the calculations of $np$ scattering and the $M1$ transition cross section of the reaction $p(n,\gamma)d$ DuWeijieHORSE  as well as to calculations of continuum nuclear spectra within various cluster models with phenomenological interactions in Refs. monopoleC12 ; Lurie1 ; Lurie2 ; Lurie3 ; Lurie4 and within the RGM 6He6He ; VVasil ; Lashko . For many-body ab initio applications, however, the HORSE method is impractical due to the prohibitive computational cost. To get around this, a simplified HORSE version, the single-state HORSE (SS-HORSE) method has been developed AMShirokovHORSE2016 ; BlokhintsevSSHorseC ; AMShirokovHORSE2018 , which makes it possible to extend the NCSM to calculations of phase shifts. Moreover, the SS-HORSE method can also provide resonance energies and widths via $S$-matrix pole location. The SS-HORSE method in combination with the NCSM was used to describe resonances in various light nuclei, in particular, in 5He, 5Li AMShirokovHORSE2016 ; AMShirokovHORSE2018 , 7He PPN ; Igor2022 , 9Li 9Li , to propose the first description of the tetraneutron resonance AMShirokov2016 ; 4nAIP , and to predict resonant states in the trineutron 3narXiv . The price paid for the SS-HORSE simplification is the loss of information on the continuum wave function. Without the wave functions, one is not able to perform calculations of important physical observables such as, e. g., electromagnetic dissociation cross sections.
The modified Hulthén–Kohn method extending the NCSM to the continuum spectrum was suggested by V. D. Efros VDEfros to address the shortcomings of the HORSE and SS-HORSE approaches in a computationally feasible manner while retaining the predictive power of the ab initio nature of the NCSM. Moreover, unlike the SS-HORSE method, the Efros method easily extends to multichannel problems. To date, this approach has been tested using simple model problems only in Ref. VDEfros .
The aim of this work is to explore in detail the possibilities of the Efros method VDEfros in solving scattering problems with the goal of utilizing it in future many-body applications. We employ this improved formalism to calculate and analyze phase shifts and continuum wave function convergence for both single and coupled channels using the Daejeon16 $NN$ interaction Daejeon16  and model two-body problems. We demonstrate the utilization of the Efros method for locating $S$-matrix poles to find resonance energies $E_{r}$ and widths $\Gamma$ as well as to improve variational results for the bound state energies $E_{b}$. We combine the ideas of the Efros method with some achievements of the HORSE formalism BangHORSE ; Zaitsev1998 which is a particular version of the $J$-matrix formalism Yamani  in quantum scattering theory. The $J$-matrix utilizes not only the oscillator basis expansions but also those in terms of the Laguerre functions which make it possible to include naturally the Coulomb interactions in the calculations. For many-body nuclear applications, however, the basis of Laguerre functions is not as convenient as the harmonic oscillator basis and we therefore focus here on the HORSE version of the $J$-matrix. As a result, we obtain both continuum and bound state wave functions as infinite series of oscillator functions. We discuss here only the scattering of uncharged particles; various extensions of the HORSE formalism accounting for the Coulomb distortion of wave function asymptotics can be found in Refs. Okhrimenko ; BangHORSE .
In Section II, we introduce the formalism of the Efros method. We discuss applications of this method to single- and coupled-channel $np$ scattering in Sections III and IV, respectively. Calculations of $S$-matrix poles are examined in Section V. We conclude in Section VI.
II Formalism
We start with presenting the Efros method VDEfros  in the case of a two-body problem with $w$ open channels and some modifications and extensions that we suppose will be useful for many-body applications.
II.1 Ansatz wave function
The scattering wave function at the energy $E$ in the center-of-mass system corresponding to the entrance channel $i$ is parameterized as
$u_{i}(E;\vec{r})=\eta^{-}_{i}(\vec{r})-\sum_{j=1}^{w}\eta^{+}_{j}(\vec{r})\,\mathbb{S}_{ji}(E)+\sum_{q=0}^{v-1}b_{q}^{i}(E)\,\beta_{q}(\vec{r}).$ (1)
Here $\vec{r}$ is the coordinate of relative separation of colliding fragments, the channel multi-indices $i$ and $j$ include all quantum numbers distinguishing the channels and those common for all coupled channels like the orbital momentum of relative motion of colliding particles $l_{j}$ in the channel $j$, their total spin $S_{j}$, and the total angular momentum $J$ and its projection $M_{J}$. $\beta_{q}(\vec{r})$ are $v$ linearly-independent short-range functions (SRFs) vanishing as ${r\rightarrow\infty}$ which are introduced to describe the wave function at short distances. The set of $v$ energy-dependent coefficients $b_{q}^{i}(E)$, $q=0,...\>,v-1$, together with $w$ $S$-matrix elements $\mathbb{S}_{ji}(E)$, $j=1,...\>,w$, comprise a set of $N=v+w$ unknowns in each given entrance channel $i$ that will be determined within this approach.
For nuclear physics applications, it is natural and convenient to use the expansions in terms of harmonic oscillator functions; therefore we set SRFs $\beta_{q}(\vec{r})$ to be equal to finite sums of these functions,
$\beta_{q}(\vec{r})=\sum_{j=1}^{w}\sum_{n=0}^{M_{j}-1}a_{qn}^{j}\phi_{nj}(\vec{r}),$ (2)
or just to single oscillator functions $\phi_{nj}(\vec{r})$ in different channels $j$ supposing that all coefficients $a_{qn}^{j}$ in Eq. (2) with an exception of one of them for each value of $q$ are equal to zero. The oscillator functions $\phi_{nj}(\vec{r})$ are defined as
$\phi_{nj}(\vec{r})\equiv\phi_{nl_{j}}\!(\vec{r})=R_{nl_{j}}\!(r)\,\mathscr{Y}_{JM_{J}}^{l_{j}S_{j}}(\Omega_{\hat{r}}),$ (3)
where $\mathscr{Y}_{JM_{J}}^{l_{j}S_{j}}(\Omega_{\hat{r}})$ are generalized spherical harmonics which describe the spin-angular degrees of freedom in the channel $j$,
$\mathscr{Y}_{JM_{J}}^{lS}(\Omega_{\hat{r}})=\sum_{m_{S},m_{l}}(lm_{l}Sm_{S}|JM_{J})Y_{lm_{l}}(\Omega_{\hat{r}})\chi_{m_{S}}^{S},$ (4)
$(lm_{l}Sm_{S}|JM_{J})$ are real-valued Clebsch–Gordan coefficients Varsha , $\chi_{m_{S}}^{S}$ is the spinor, and $Y_{lm_{l}}(\Omega_{\hat{r}})$ is the spherical harmonic function Varsha ;
$R_{nl}(r)=(-1)^{n}\sqrt{\frac{2n!}{b^{3}\Gamma(n+l+\frac{3}{2})}}\Big{(}\frac{r}{b}\Big{)}^{l}e^{-\frac{r^{2}}{2b^{2}}}L_{n}^{l+\frac{1}{2}}\!\Big{(}\frac{r^{2}}{b^{2}}\Big{)},$ (5)
where $L_{n}^{\alpha}(z)$ is the associated Laguerre polynomial Arfken ,
$b=\sqrt{\frac{\hbar}{m\Omega}}$ (6)
is the oscillator length parameter, $m$ is the reduced mass of the colliding particles, and $\Omega$ is the oscillator frequency.
The long-range behavior of the wave functions $u_{i}(E;\vec{r})$ is described by the functions
$\eta^{\pm}_{i}(\vec{r})\equiv\eta^{\pm}_{i}(k,\vec{r})=\tilde{h}^{\pm}_{l_{i}}(kr)\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}}).$ (7)
Here
$\tilde{h}^{\pm}_{l}(kr)=\tilde{n}_{l}(kr)\pm ij_{l}(kr),$ (8)
$k=\frac{\sqrt{2mE}}{\hbar}$ (9)
is the relative momentum of colliding particles, $j_{l}(z)$ is the spherical Bessel function Arfken , and the functions $\tilde{n}_{l}(z)$ and $\tilde{h}^{\pm}_{l}(z)$ behave asymptotically as
$\tilde{n}_{l}(z)$ $\mathop{\longrightarrow}_{z\to\infty}-n_{l}(z)\mathop{\longrightarrow}_{z\to\infty}\frac{1}{z}\cos\!\Big{(}z-\frac{1}{2}\pi l\Big{)},$ (10)
$\tilde{h}^{\pm}_{l}(z)$ $\mathop{\longrightarrow}_{z\to\infty}{\pm ih^{(1,2)}_{l}(z)}\mathop{\longrightarrow}_{z\to\infty}\frac{i^{\mp l}}{z}e^{\pm iz}.$ (11)
The irregularity at the origin of the spherical Neumann $n_{l}(z)$ and Hankel $h^{(1,2)}_{l}(z)$ functions Arfken  should be regularized to avoid divergences of integrals calculated with them.
In the case of the continuum spectrum, the series of oscillator functions $\phi_{nj}(\vec{r})$ should be formally infinite:
$u_{i}(E;\vec{r})=\sum_{j=1}^{w}\sum_{n=0}^{\infty}d_{n,ij}(E)\,\phi_{n{j}}(\vec{r}).$ (12)
Therefore, contrary to Efros VDEfros  who uses regulator functions for the regularization of the functions $n_{l}(z)$ and $h^{(1,2)}_{l}(z)$, we utilize here the functions developed within the HORSE formalism BangHORSE ; Zaitsev1998 defined through their infinite expansions in a series of harmonic oscillator functions both for the regularized functions $\tilde{n}_{l}(kr)$ and $\tilde{h}^{\pm}_{l}(kr)$ and for the regular spherical Bessel function $j_{l}(kr)$:
$j_{l}(kr)=$ $\sum_{n=0}^{\infty}S_{nl}(k)R_{nl}(r),$ (13)
$\tilde{n}_{l}(kr)=$ $\sum_{n=0}^{\infty}C_{nl}(k)R_{nl}(r),$ (14)
$\tilde{h}^{\pm}_{l}(kr)=$ $\sum_{n=0}^{\infty}C^{\pm}_{nl}(k)R_{nl}(r),$ (15)
where BangHORSE
$S_{nl}(k)=$ $\>\sqrt{\frac{\pi b^{3}n!}{\Gamma(n+l+\frac{3}{2})}}(kb)^{l}e^{-\frac{k^{2}b^{2}}{2}}L_{n}^{l+\frac{1}{2}}(k^{2}b^{2}),$ (16)
$C_{nl}(k)=$ $\>\frac{(-1)^{l}}{\Gamma(-l+\frac{1}{2})}\sqrt{\frac{\pi b^{3}n!}{\Gamma(n+l+\frac{3}{2})}}(kb)^{-l-1}e^{-\frac{k^{2}b^{2}}{2}}$
×$\phantom{\Gamma(-l)}\times{{}_{1}F_{1}}\Big{(}\!-n-l-\frac{1}{2},-l+\frac{1}{2};k^{2}b^{2}\Big{)},$ (17)
$C^{\pm}_{nl}(k)=$ $\>C_{nl}(k)\pm iS_{nl}(k),$ (18)
and ${{}_{1}F_{1}}(a,b;z)$ is the confluent hypergeometric function Arfken .
It is convenient sometimes to employ the standing-wave approach where the continuum wave functions are real-valued:
${u}_{i}(E;\vec{r})=j_{l_{i}}(kr)\,\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})\\ +\sum_{j=1}^{w}K_{ji}(E)\,\tilde{n}_{l_{j}}(kr)\,\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})+\sum_{q=0}^{v-1}{b}_{q}^{i}(E)\,\beta_{q}(\vec{r}),$ (19)
where we define the $K$-matrix as
$K=i(I-\mathbb{S})(I+\mathbb{S})^{-1}\!.$ (20)
II.2 Formal solutions
The wave functions $u_{i}(E;\vec{r})$ entering Eq. (1) fit the Schrödinger equation
$(H-E)\,u_{i}(E;\vec{r})=0,$ (21)
where $H$ is the Hamiltonian of the system. To find $N$ unknowns entering the function $u_{i}(E;\vec{r})$ in each entrance channel $i$, we can obtain a set of $w$ matrix equations
[…]
$\bar{\beta}_{q}^{i}(\vec{r})=\sum_{j=1}^{w}\sum_{n=0}^{\bar{M}_{j}-1}\bar{a}_{qn;i}^{j}\phi_{nj}(\vec{r}).$ (23)
Efros suggests VDEfros to use the same $v$ SRFs $\bar{\beta}_{q}(\vec{r})=\beta_{q}(\vec{r})$, $q=0,...\,,v-1,$ and we shall mostly follow this suggestion in what follows. In this case, $\bar{a}_{qn;i}^{j}=a_{qn}^{j}$ for $q=0,...\,,{v-1}$. For $q=v,...\,,N$, $\bar{\beta}_{q}(\vec{r})$ may be arbitrarily chosen with the only constraint being that the functions $\bar{\beta}_{q}(\vec{r}),q=0,...,N$ should be linearly-independent. One of the reasonable choices, for example, may be some oscillator functions as we shall see below.
However, generally, the set of SRFs $\{\bar{\beta}_{q}\}$ can differ completely from the set of linearly-independent SRFs $\{{\beta}_{q}\}$; it is only important to have a set of $N$ linearly-independent SRFs $\bar{\beta}_{q}(\vec{r})$ which may be non-orthogonal. That means that, generally, the coefficients $\bar{a}_{qn;i}^{j}$ are distinct from $a_{qn}^{j}$ in Eq. (2). Moreover, $\bar{M}_{j}$, the finite truncations of oscillator series in each channel $j$, may differ from the truncations $M_{j}$ in Eq. (2).
Let us define a vector $\bar{c}^{i}$ of the unknowns in the entrance channel $i$ with components
$\bar{c}^{i}_{q}=\begin{cases}\phantom{-}b_{q}^{i}(E),\quad&q=0,...\,,v-1,\\ -\mathbb{S}_{q-v+1,i}(E),&q=v,...\,,N-1.\end{cases}$ (24)
Equation (22) can be rewritten as
$\bar{A}\bar{c}^{i}=\bar{B}^{i},\quad i=1,...\,,w,$ (25)
where the elements of matrix $\bar{A}$ are
$\bar{A}_{qq^{\prime}}=\langle{\bar{\beta}_{q}}|H-E|\beta_{q^{\prime}}\rangle,\quad$$\begin{cases}q=0,...\,,N-1,\\ q^{\prime}=0,...\,,v-1,\end{cases}$ (26a)
$\bar{A}_{qq^{\prime}}={\langle{\bar{\beta}_{q}}|H-E|\eta^{+}_{{j}}\rangle},\quad$$\begin{cases}{q=0,...\,,N-1},\\ q^{\prime}=v,...\,,N-1,\\ {j=q^{\prime}-v+1,}\end{cases}$ (26b)
and the components of vector $\bar{B}^{i}$ are
$\bar{B}_{q}^{i}=-\langle{\bar{\beta}_{q}}|H-E|\eta^{-}_{{i}}\rangle,\quad$ ${q=0,...\,,N-1}.$ (26c)
We can combine $w$ equations (25) corresponding to different entrance channels $i$ into a single matrix equation
$Ac=B,$ (27)
where
$A=\begin{pmatrix}\bar{A}&\textbf{0}&\ldots&\textbf{0}\\[3.00003pt] \textbf{0}&\bar{A}&\ldots&\textbf{0}\\[3.00003pt] \vdots&\vdots&\ddots&\vdots\\[3.00003pt] \textbf{0}&\textbf{0}&\ldots&\bar{A}\\[3.00003pt] \end{pmatrix}$ (28)
is a $wN\times wN$ matrix comprising $w$ identical $N\times N$ matrices $\bar{A}$ as blocks on its diagonal, and the $wN$-dimensional vectors $B$ and $c$ are
$B=\begin{pmatrix}\bar{B}^{1}\\[3.00003pt] \vdots\\[3.00003pt] \bar{B}^{w}\\[3.00003pt] \end{pmatrix}\!$ (29)
and
$c=\begin{pmatrix}\bar{c}^{1}\\[3.00003pt] \vdots\\[3.00003pt] \bar{c}^{w}\\[3.00003pt] \end{pmatrix}\!.$ (30)
By solving Eq. (27) we can obtain all $wN$ unknowns entering Eqs. (1) with different entrance channels $i=1,...\,,w$. In the single-channel case, the above equations provide a complete solution of the problem. However, not all the $Nw$ unknowns are independent in the multichannel case when $w>1$. In particular, the $S$ matrix should be symmetric, i. e., $\mathbb{S}_{ij}(E)=\mathbb{S}_{ji}(E)$. So, the $S$ matrix has $w^{2}$ matrix elements but only ${\frac{1}{2}w(w+1)}$ independent matrix elements. Therefore, Eq. (27) is over-complete, providing ${w^{2}-\frac{1}{2}w(w+1)}=\frac{1}{2}w(w-1)$ unneeded extra relations between unknowns, and the number of independent unknowns is ${\mathbb{\color[rgb]{0,0,0}N}}{=Nw-\frac{1}{2}w(w-1)}$. We shall discuss this situation below using an example of a two-channel system. We shall show that in some cases the over-complete Eq. (27) provides a correct solution with $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$. However, in cases of a reduced set of SRFs $\bar{\beta}_{q}$ most interesting for many-body applications, solving Eq. (27) results in $\mathbb{S}_{12}(E)\neq\mathbb{S}_{21}(E)$.
In the cases where the symmetry is not guaranteed, we impose the $S$-matrix symmetry and replace all $S$-matrix elements $\mathbb{S}_{ij}(E)$ with $i<j$ by $\mathbb{S}_{ji}(E)$ in Eqs. (1) and (22) and define a new vector of unknowns $c$ by replacing the vectors $\bar{c}^{i}$ defined by Eq. (24) by the vectors $c^{i}$ with the components
${c}^{i}_{q}=\begin{cases}\phantom{-}b_{q}^{i}(E),\quad&q=0,...\,,v-1,\\ -\mathbb{S}_{q-v+i,i}(E),&q=v,...\,,N-i.\end{cases}$ (31)
We combine the vectors ${c}^{i}_{q}$ in a total ${\mathbb{\color[rgb]{0,0,0}N}}$-dimensional vector of unknowns
$c=\begin{pmatrix}c^{1}\\[3.00003pt] \vdots\\[3.00003pt] c^{w}\\[3.00003pt] \end{pmatrix}\!.$ (32)
Instead of $wN$ equations (22) we need now to formulate only $\mathbb{\color[rgb]{0,0,0}N}$ independent equations of the same type determining the $\mathbb{\color[rgb]{0,0,0}N}$ independent unknowns. We can do this by reducing the sets of bra SRFs $\langle{\bar{\beta}_{q}}|$ in Eq. (22) in some entrance channels $i$. We denote these reduced sets as $\{\langle\bar{\beta}_{q}^{i}|\}$. We include in the set $\{\langle\bar{\beta}_{q}^{1}|\}$ in the first entrance channel ($i=1$) all $N$ SRFs $\langle{\bar{\beta}_{q}}|$, $\{\langle\bar{\beta}_{q}^{1}|\}\equiv\{\langle{\bar{\beta}_{q}}|\}$. We note here that the channel ordering is arbitrary and we can always choose a particular numbering of channels which is optimal from the point of view of providing the best convergence. For the second entrance channel with $i=2$, we use $N-1$ independent SRFs $\langle\bar{\beta}_{q}^{2}|$ which may be obtained from the set $\{\langle{\bar{\beta}_{q}}|\}$ by excluding from it one of the SRFs $\langle{\bar{\beta}_{q}}|$. We can also define some of the SRFs $\langle\bar{\beta}_{q}^{2}|$ as linear combinations of the SRFs $\langle{\bar{\beta}_{q}}|$: the set $\{\langle\bar{\beta}_{q}^{2}|\}$ should include $N-1$ independent SRFs which can be non-orthogonal. We note that any superposition of $v$ linearly-independent SRFs $\beta_{q}(\vec{r})$, $q=0,...\,,v-1$, entering Eq. (1) should produce exactly the same results while superpositions involving additional $w$ SRFs ${\bar{\beta}_{q}(\vec{r})}$, $q=v,...\,,v+w$, involved in Eq. (22) may affect the convergence. Next, we obtain in the same manner the set $\{\langle\bar{\beta}_{q}^{3}|\}$ including $N-2$ independent SRFs $\langle\bar{\beta}_{q}^{3}|$, etc. We note that the set of ket SRFs $\{|\beta_{q^{\prime}}\rangle\}$ is the same in all entrance channels.
As a result, instead of Eq. (25), we obtain
${A}^{i}c^{i}=B^{i},\quad i=1,...\,,w,$ (33)
where the matrices ${A}^{i}$, contrary to matrices $\bar{A}$ in Eq. (25), differ in size for different channels $i$ and their elements are
$\!\!{A}^{i}_{qq^{\prime}}$ $=\langle\bar{\beta}^{i}_{q}|H-E|\beta_{q^{\prime}}\rangle,\;$$\begin{cases}q=0,...\,,N-i,\\ q^{\prime}=0,...\,,v-1,\end{cases}$ (34a)
$\!{A}^{i}_{qq^{\prime}}$ $=\langle\bar{\beta}^{i}_{q}|H-E|\eta^{+}_{{j}}\rangle,\;$$\begin{cases}{q=0,...\,,N-i},\\ q^{\prime}=v,...,N-i,\color[rgb]{0,0,0}\\ {j=q^{\prime}-v+i.}\color[rgb]{0,0,0}\end{cases}$ (34b)
For the components of vector $B^{i}$ we define for $i=1$ as
$B^{1}_{q}=-\langle\bar{\beta}^{1}_{q}|H-E|\eta^{-}_{{i}}\rangle,\;\qquad$ $q=0,...\,,N-1,$ (34c)
and for $i>1$ as
$B_{q}^{i}=-\langle\bar{\beta}^{i}_{q}|H-E|{\eta^{-}_{{i}}}\rangle-\sum_{j=1}^{i-1}$ $\langle\bar{\beta}^{i}_{q}|H-E|\eta^{+}_{{j\color[rgb]{0,0,0}}}\rangle c^{j}_{v+i-j\color[rgb]{0,0,0}},$
$q=0,...\,,N-i.$ (34d)
Equations (33) can be combined into the matrix equation (27) where
$A=\begin{pmatrix}{A}^{1}&\textbf{0}&\ldots&\textbf{0}\\[3.00003pt] \textbf{0}&{A}^{2}&\ldots&\textbf{0}\\[3.00003pt] \vdots&\vdots&\ddots&\vdots\\[3.00003pt] \textbf{0}&\textbf{0}&\ldots&{A}^{w}\\[3.00003pt] \end{pmatrix}$ (35)
is an $\mathbb{\color[rgb]{0,0,0}N}\times\mathbb{\color[rgb]{0,0,0}N}$ matrix comprising $(N-i+1)\times(N-i+1)$ matrices $A^{i}$ as diagonal blocks,
$B=\begin{pmatrix}B^{1}\\[3.00003pt] \vdots\\[3.00003pt] B^{w}\\[3.00003pt] \end{pmatrix}\!,$ (36)
and $c$ given by Eq. (32) are $\mathbb{\color[rgb]{0,0,0}N}$-dimensional vectors.
Equations (33) can be resolved step-by-step. First, we solve this matrix equation for the first entrance channel with $i=1$ and obtain $v$ coefficients $b_{q}^{1}(E)={c}^{1}_{q}$, $q=0,...\,,v-1$, and $w$ $S$-matrix elements ${\mathbb{S}_{j1}(E)=-c^{1}_{j+v-1}}$, $j=1,...\,,w$ at any desired energy $E$. Next, we solve Eq. (33) for $i=2$. The $S$-matrix element ${\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)=-c^{1}_{v+1}}$ in the right-hand-side of Eq. (34d) was already obtained at the previous step. With the help of Eq. (33) we now obtain $v$ coefficients ${b_{q}^{2}(E)={c}^{2}_{q}}$, $q=0,...\,,v-1$, and $w-1$ $S$-matrix elements ${\mathbb{S}_{j2}(E)=-c^{2}_{j+v-2}}$, $j=2,...\,,w$. Now we have $S$-matrix elements $\mathbb{S}_{3j}(E)=-c^{j}_{v+3-j}$ with $j=1$, 2 in the right-hand-side of Eq. (33) for $i=3$ and solve it. Next we solve Eq. (33) for $i=4$, 5, etc.
The calculations of matrices $A^{i}$ and $B^{i}$ can be simplified using the following expressions for the function $\eta^{\pm}_{i}(\vec{r})$ Zaitsev1998 :
$(T-E)\eta^{\pm}_{i}(\vec{r})=\frac{\hbar^{2}}{2m}\frac{\phi_{0i}(\vec{r})}{k\,S_{0l_{i}}(k)},$ (37)
where $T$ is the operator of kinetic energy of relative motion of colliding particles. The Hamiltonian
$H=T+V$ (38)
includes also the potential energy operator $V$. Thus,
$(H-E)\eta^{\pm}_{i}(\vec{r})=V\eta^{\pm}_{i}(\vec{r})+\frac{\hbar^{2}}{2m}\frac{\phi_{0i}(\vec{r})}{k\,S_{0l_{i}}(k)}.$ (39)
Following the ideas of the HORSE approach BangHORSE ; Zaitsev1998 , the potential energy operator $V$ can be approximated by a finite matrix $V^{tr}$ in the oscillator basis,
$V\approx V^{tr}=\sum_{i,j=1}^{w}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle\langle\phi_{ni}|V|\phi_{n^{\prime}j}\rangle\langle\phi_{n^{\prime}j}|\\ \equiv\sum_{i,j=1}^{w}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle V_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|,$ (40)
since the potential energy matrix elements in the oscillator basis $V_{ni,n^{\prime}j}$ decrease with $n$ and/or $n^{\prime}$ and can be neglected at large $n$ and/or $n^{\prime}$ as compared to the non-zero kinetic energy matrix elements $T_{ni,n^{\prime}j}=\langle\phi_{ni}|T|\phi_{n^{\prime}j}\rangle=T^{l_{i}}_{nn^{\prime}}\delta_{ij}$,
$T^{l}_{nn}$ $=\frac{1}{2}\hbar\Omega\Big{(}2n+l+\frac{3}{2}\Big{)},$ (41a)
$T^{l}_{n-1,n}$ $=-\frac{1}{2}\hbar\Omega\sqrt{n\Big{(}n+l+\frac{1}{2}\Big{)}},$ (41b)
$T^{l}_{n+1,n}$ $=-\frac{1}{2}\hbar\Omega\sqrt{(n+1)\Big{(}n+l+\frac{3}{2}\Big{)}},$ (41c)
$T_{nn^{\prime}}^{l}$ $=0\quad\text{if}\quad{\color[rgb]{0,0,0}|n-n^{\prime}|>1.}$ (41d)
Note, the approximation of Eq. (40) is conventionally employed in many-body ab initio nuclear structure calculations where the kinetic energy matrix (41) is usually also truncated in the same manner. However, we do not truncate the kinetic energy to allow for scattering.
Using Eq. (37) together with Eqs. (2), (7), (15), (23), and defining
$H_{ni,n^{\prime}j}=\langle\phi_{ni}|H|\phi_{n^{\prime}j}\rangle,$ (42)
we rewrite Eqs. (34) as
$\!\!\!{A}^{i}_{qq^{\prime}}=$ ${\sum_{j,j^{\prime}=1}^{w}}\sum_{n=0}^{\bar{M}_{j}-1}\sum_{n^{\prime}=0}^{M_{j^{\prime}}-1}{\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}}a_{q^{\prime}n^{\prime}}^{j^{\prime}}(H_{nj,n^{\prime}j^{\prime}}-E\delta_{nn^{\prime}}\delta_{jj^{\prime}}),\quad$$\begin{cases}q=0,...\,,N-i,\\ q^{\prime}=0,...\,,v-1,\end{cases}$ (43a)
$\!\!\!{A}^{i}_{qq^{\prime}}=$ $\sum_{j=1}^{w}\sum_{n=0}^{\bar{\cal M}_{j}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j^{\prime}}-1}\color[rgb]{0,0,0}\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}C^{+}_{n^{\prime}l_{\!j^{\prime}}}\!(k)\,V_{nj,n^{\prime}j^{\prime}}+\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{k\,S_{0l_{\!j^{\prime}}}\!(k)},\quad$$\begin{cases}{q=0,...\,,N-i},\\ q^{\prime}=v,...,N-i,\\ {j^{\prime}=q^{\prime}-v+i,}\end{cases}\!\!$ (43b)
$\!\!B^{1}_{q}=$ $-\sum_{j=1}^{w}\sum_{n=0}^{\bar{\cal M}_{j}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{1}-1}\color[rgb]{0,0,0}\big{(}\bar{a}_{qn;1}^{j}\big{)}^{\!*}C^{-}_{n^{\prime}l_{1}}\!(k)\,V_{nj,n^{\prime}1}-\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;1}^{1}\big{)}^{\!*}}{k\,S_{0l_{1}}\!(k)},\quad$ $q=0,...\,,N-1,\!$ (43c)
$\!\!B_{q}^{i}=$ $-\sum_{j=1}^{w}\sum_{n=0}^{\bar{\cal M}_{j}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{i}-1}\color[rgb]{0,0,0}\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}C^{-}_{n^{\prime}l_{i}}\!(k)\,V_{nj,n^{\prime}i}-\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{i}\big{)}^{\!*}}{k\,S_{0l_{i}}\!(k)}$
$-\sum_{j^{\prime}=1}^{i-1}\sum_{j=1}^{w}\sum_{n=0}^{\bar{\cal M}_{j}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j^{\prime}}-1}\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}C^{+}_{n^{\prime}l_{j^{\prime}}}\!(k)\,V_{nj,n^{\prime}j^{\prime}}c^{j^{\prime}}_{v+i-j^{\prime}}-\sum_{j^{\prime}=1}^{i-1}\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{kS_{0l_{j^{\prime}}}(k)}c^{j^{\prime}}_{v+i-j^{\prime}},\quad$$\begin{cases}i>1,\\ q=0,...\,,N-i,\end{cases}$ (43d)
where ¯$\bar{\mathcal{M}}_{i}=\min\{\bar{M}_{i},\mathcal{N}_{i}\}$, $\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}$ is a complex conjugate of $\bar{a}_{qn;i}^{j}$, and, as previously mentioned, we set $\bar{a}_{qn;i}^{j}=a_{qn}^{j}$ for $q=0,...,{v-1}$ and arbitrarily chosen coefficients for $q=v,...,{N-i}$.
Efros mentioned VDEfros that, contrary to the conventional Hulthén–Kohn variational method Kohn1948 ; Moses1953 ; Moses , his approach does not make use of “free-free” matrix elements like $\langle\eta^{\pm}_{i}|H-E|\eta^{\pm}_{j}\rangle$ but, instead, includes only “bound-bound” $\langle\bar{\beta}^{i}_{q}|{H-E}|\beta_{q^{\prime}}\rangle$ and “bound-free” $\langle\bar{\beta}^{i}_{q}|{H-E}|\eta^{\pm}_{j}\rangle$ matrix elements which are easier to calculate. We note that Eqs. (43) include matrix elements of the “bound-bound” type only, which are evaluated within a restricted Hilbert space spanned by a limited set of oscillator functions.
We should also note that in many-body applications, Eqs. (43b)–(43d) become more complicated. The problem is that the terms ±$\pm\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{k\,S_{0l_{j^{\prime}}}\!(k)}$ in these equations account for the kinetic energy of relative motion of colliding particles $T^{rel}$ only. In the many-body case, the colliding particles have an internal structure. Therefore, the Hamiltonian includes the kinetic energy of the internal motion of these particles $T^{int}$, which enters Eqs. (43b)–(43d) in the same manner as potential energy terms. The separation of the total kinetic energy operator $T=T^{rel}+T^{int}$ into the relative $T^{rel}$ and internal $T^{int}$ parts may be non-trivial within NCSM but will be much more natural and simpler in approaches like the RGM.
In the case of the standing-wave approach starting from the wave-function ansatz (19), one can use all the above equations of this subsection with replacements $-\mathbb{S}_{ij}(E)\rightarrow K_{ij}(E)$, $\tilde{h}^{-}_{l_{i}}(kr)\rightarrow j_{l_{i}}(kr)$, ${\tilde{h}^{+}_{l_{i}}(kr)\rightarrow\tilde{n}_{l_{i}}(kr)}$, $C^{-}_{nl_{i}}(k)\rightarrow S_{nl_{i}}(k)$, ${C^{+}_{nl_{i}}(k)\rightarrow C_{nl_{i}}(k)}$, having in mind that with these replacements, Eqs. (37) should be substituted by the following relations:
$(T-E)j_{l_{i}}(kr)\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})$ $=0,$ (37a)
$(T-E)\tilde{n}_{l_{i}}(kr)\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})$ $=\frac{\hbar^{2}}{2m}\frac{\phi_{ni}(\vec{r})}{k\,S_{0l_{i}}(k)}.$ (37b)
As a result, Eqs. (43a)–(43b) preserve their structure with the replacement $C^{+}_{nl}(k)\rightarrow C_{nl}(k)$ while in Eqs. (43c)–(43d) the substitution ${C^{-}_{nl}(k)\rightarrow S_{nl}(k)}$ should be accompanied by removing the terms $-\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j}\big{)}^{\!*}}{kS_{0l_{j}}\!(k)}$.
II.3 $S$-matrix poles
The matrix formulation of the problem presented by Eqs. (27), (32), (35), (36) is convenient for searching for the $S$-matrix poles. By locating the $S$-matrix poles in the complex energy plane, it is possible to obtain resonance energies $E_{r}$ and widths $\Gamma$ AMShirokov2016 ; 4nAIP ; Lurie3 ; Lurie4 ; AMShirokovHORSE2016 ; AMShirokovHORSE2018 ; PPN ; Igor2022 ; 9Li ; 3narXiv and to improve essentially the results of variational calculations of bound state energies Vasya21 ; Lurie3 ; Lurie4 ; 9Li ; moreover, by calculating the residues for poles corresponding to bound states, it is possible to obtain Vasya21 ; 9Li the asymptotic normalization constants Baz .
The formal solution of Eq. (27) for the vector of unknowns $c$ can be written as
$c=A^{-1}B.$ (44)
Since $c$ contains the $S$-matrix elements, it is natural to search for the $S$-matrix poles by solving the equation
$\det A(E)=0$ (45)
in the complex energy plane. Due to the diagonal structure of matrix $A$, see Eq. (35), it looks like that the solutions of Eq. (45) can be assured by solving any of the equations
$\det A^{i}(E)=0;$
however, it is safe to use solutions only of the equation
$\det A^{1}(E)=0.$ (46)
Indeed, Eq. (44) can be split into a set of equations for pieces of the vector of unknowns $c$:
$c^{i}=(A^{i})^{-1}B^{i}.$ (47)
The vector $B^{i}$ at $i>1$ includes $S$-matrix elements [the components $c^{j}_{v+i-j}$ of the vector $c$, see Eqs. (34d) and (31)] which have poles; thus searching for poles of the vector $c^{i}$ by locating the zeroes of $\det A^{i}$ with $i>1$ is not justified and one can omit some of the $S$-matrix poles. Locating the zeroes of $\det A^{1}(E)$ in the complex energy plane can be efficiently performed by the technique of calculating contour integrals described in Ref. AMShirokovHORSE2018 .
[…]
$A^{1}c^{1}=0.$ (48)
Note, Eq. (46) guarantees that Eq. (48) has a non-zero solution. However, the vector of unknowns $c^{1}$ is obtained up to an arbitrary multiplier, and the bound state wave function $u_{i}(E;\vec{r})$ should be additionally normalized. Moreover, the bound state wave function in this case is expressed as an infinite converging series of oscillator functions since the function $\eta^{+}_{i}(\vec{r})$ entering $u_{i}(E;\vec{r})$ is defined through infinite oscillator function series, see Eqs. (7) and (15).
III Single-Channel Scattering
We proceed to discuss the efficiency of the Efros approach and its advantages and disadvantages using two-body single-channel scattering. As the first example, we consider $np$ scattering in the ${{}^{1}S_{0}}$ channel using the Daejeon16 $NN$ interaction Daejeon16 . In the single-channel case, it is convenient to utilize the standing-wave approach with real-valued wave functions where the only $K$-matrix element in Eq. (19) is expressed through the scattering phase shifts $\delta_{l}(E)$:
$K_{11}=\tan\delta_{l}(E).$ (49)
We shall omit the channel index $i$ in the matrix $A^{i}$, vectors $B^{i}$ and $c^{i}$, coefficients $b_{q}^{i}(E)$, etc., while discussing single-channel applications.
III.1 Naive application
Various nuclear structure approaches, e. g., the NCSM, utilize wave function expansions in the finite series of many-body harmonic oscillator functions ${\phi_{n}(\vec{r})\equiv\phi_{nj}(\vec{r})}$ like Eqs. (2) and (3), where we exclude in our case of two-body single-channel scattering the summation over channels and can omit now the channel index $j$. In this case, the Hamiltonian has a structure
$H^{tr}=\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle\langle\phi_{n}|H|\phi_{n^{\prime}}\rangle\langle\phi_{n^{\prime}}|\equiv\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle H_{nn^{\prime}}\langle\phi_{n^{\prime}}|,$ (50)
where $\mathcal{\color[rgb]{0,0,0}N}$ is the number of harmonic oscillator basis functions used in the truncated wave function expansion. The truncation corresponds conventionally to the maximal value of (excitation) oscillator quanta $N_{\max}$. Having in mind future many-body applications combining the NCSM and the Efros method, we start with a Hamiltonian of this type naively supposing that solutions are obtainable within the Efros approach.
In the case of single-channel two-body scattering, the relative motion oscillator functions $\phi_{n}(\vec{r})$ are given by Eq. (3) and $2n+l\leq N_{\textrm{max}}$, where $n=0,1,...\,,\mathcal{\color[rgb]{0,0,0}N}-1$. Equation (12) takes the form
$u(E;\vec{r})=\sum_{n=0}^{\infty}d_{n}(E)\phi_{n}(\vec{r}),$ (51)
where the expansion coefficients are expected to be
$d_{n}(E)$ =$=\begin{cases}g_{n}(E)+a^{as}_{n}(E),&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}-1,\\ a^{as}_{n}(E),&n=\mathcal{\color[rgb]{0,0,0}N},...\,,\infty,\end{cases}$ (52)
$a^{as}_{n}(E)$ $=S_{nl}(k)+C_{nl}(k)\tan\delta_{l}(E),\quad n=0,...\,,\infty,$ (53)
and $\mathcal{\color[rgb]{0,0,0}N}$ energy-dependent coefficients $g_{n}(E)$ should be found within the approach. In our case of $NN$ scattering in the ${}^{1}S_{0}$ partial wave, the fixed quantum numbers are $l=0$, $S=0$, $J=0$, and $M_{J}=0$; therefore ${\mathcal{\color[rgb]{0,0,0}N}=\frac{1}{2}N_{\max}+1}$.
We need to specify the sets of SRFs $\{\beta_{q}\}$ and ${\{\bar{\beta}_{q}\}\equiv\{\bar{\beta}^{1}_{q}\}}$ entering Eqs. (34). A natural option is to use $N\leq\mathcal{\color[rgb]{0,0,0}N}$ eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ of the truncated Hamiltonian (50),
$\bar{\beta}_{q}(\vec{r})$ $=\beta_{q}^{\textrm{Efn}}(\vec{r}),\qquad q=0,...\,,N-1,$ (54a)
${\beta}_{q}(\vec{r})$ $=\beta_{q}^{\textrm{Efn}}(\vec{r}),\qquad q=0,...\,,v=N-2,$ (54b)
which solve the Schrödinger equation
$H^{tr}\beta_{q}^{\textrm{Efn}}(\vec{r})=E_{q}\beta_{q}^{\textrm{Efn}}(\vec{r})$ (55)
and can be expressed as superpositions of oscillator functions (3),
$\beta_{q}^{\textrm{Efn}}(\vec{r})=\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}a^{\textrm{Efn}}_{qn}\phi_{n}(\vec{r}).$ (56)
In this case, $\mathcal{\color[rgb]{0,0,0}N}$ coefficients $g_{n}(E)=\sum_{q=0}^{N-2}b_{q}(E)\,a^{\textrm{Efn}}_{qn}$ in Eq. (52) are expressed through ${v=N-1}$ unknown energy-dependent coefficients $b_{q}(E)$ in the wave function expansion (1) which should be found within the approach.
The use of the eigenfunctions simplifies the calculation of matrix $A=A^{1}$:
$A_{qq^{\prime}}=(E_{q}-E)\delta_{qq^{\prime}},\quad q=0,...\,,N-1;\quad q^{\prime}=0,...\,,N-2.$ (57)
The choice of eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as SRFs is also convenient from the point of view of many-body applications in combination with the NCSM or other ab initio approaches. These approaches conventionally utilize the M-scheme and thus operate with a huge number of many-body basis functions (Slater determinants) that do not have a definite angular momentum $J$. Therefore, to simplify applications of the Efros method, it is preferable to organize superpositions of these basis functions with a definite $J$ value. The available NCSM codes like MFDn MFDnMaris2010 ; Euro-Par ; Aktulga ; NCSM-NTSE2014 ; Shao and BIGSTICK BIGSTICK provide the NCSM eigenenergies and eigenfunctions with definite $J$ values obtained dynamically during the Hamiltonian diagonalization. So, the most natural and simple way to arrange definite $J$ superpositions of M-scheme Slater determinants is to use the NCSM eigenfunctions corresponding to eigenenergies in the energy range of interest.
Setting $\bar{\beta}_{q}=\beta_{q}^{\textrm{Efn}^{\vphantom{P}}}(\vec{r})$ in Eq. (22) and using the standing-wave functions ${u}(E;\vec{r})$ given by Eq. (19), with the help of Eq. (55) we obtain:
$(E_{q}-E)\langle\beta^{\textrm{Efn}}_{q}|u(E;\vec{r})\rangle=0,\quad q=0,...\,,N-1.$ (58)
Therefore, at any energy $E\neq E_{q}$,
$\langle\beta_{q}^{\textrm{Efn}}|{u}(E;\vec{r})\rangle=0,\quad q=0,...\,,N-1,$ (59)
and the last of these equations with $q=N-1$ results in
$\tan\delta_{l}(E)=\frac{B_{N-1}}{A_{N-1,N-1}}.$ (60)
The resulting phase shifts $\delta_{0}(E)$ in the ${}^{1}S_{0}$ partial wave of $np$ scattering obtained with the Daejeon16 $NN$ interaction Daejeon16 using the oscillator basis with ${\hbar\Omega=30}$ MeV truncated at $N_{{\max}}=16$ are presented in Fig. 1. It is seen that the obtained phase shifts at any $N$ have nothing to do with the exact ones. This issue stems from the structure of the Hamiltonian $H^{tr}$ in Eq. (50) and the use of its eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as SRFs $\bar{\beta}_{q}(\vec{r})$ which result in producing the wave function ${u}(E;\vec{r})$ as an expansion in a truncated set of oscillator functions $\phi_{n}(\vec{r})$ with ${n\leq\mathcal{\color[rgb]{0,0,0}N}-1}$ with no coupling to the continuum. The wave function components $\phi_{n}(\vec{r})$ with $n=\mathcal{\color[rgb]{0,0,0}N},\mathcal{\color[rgb]{0,0,0}N}+1,...\,,\,\infty$ are also needed since only the infinite sum of oscillator functions in Eq. (51) can describe the continuum spectrum wave function. Therefore, to calculate the contribution of the wave function components $\phi_{n}(\vec{r})$ with $n\geq\mathcal{\color[rgb]{0,0,0}N}$, we obtain an oversimplified Eq. (60) for $\tan\delta_{l}(E)$ which enters Eq. (53) for $a^{as}_{n}(E)$. Equation (60) contains information only about a single eigenstate $\beta_{N-1}^{\textrm{Efn}}(\vec{r})$ of the Hamiltonian $H^{tr}$ and thus cannot describe its complete structure and the interaction $V$ in the system. So, we cannot expect to obtain adequate phase shifts using Eq. (60) in a range of energies $E$.
Figure 1: ${}^{1}S_{0}$ phase shifts $\delta_{0}(E)$ as functions of energy $E$ obtained for various $N$ with $N_{\textrm{max}}=16$ and $\hbar\Omega=30$ MeV using eigenfunctions $\beta_{q}^{\textrm{Efn}}$ as SRFs. Vertical lines indicate the eigenvalues of $H^{tr}$.
One can, however, expect that in this case we obtain at least good phase shifts $\delta_{l}(E_{q})$ at energies $E=E_{q}$ [see Eq. (58)] and can use them for evaluating phase shifts in some energy interval as is done within the SS-HORSE approach AMShirokovHORSE2016 ; BlokhintsevSSHorseC ; AMShirokovHORSE2018 . We see in Fig. 1 that the curves describing phase shifts obtained with different $N$ cross the curve representing the exact phase shifts at energies close to the lowest $E_{q}$ values, however, contrary to the SS-HORSE, calculations of $\tan\delta_{l}(E_{q})$ using Eq. (60) are not accurate and this inaccuracy increases for higher eigenenergies.
Another reasonable option for SRFs $\beta_{q}$ and $\bar{\beta}_{q}$ are the oscillator functions:
$\bar{\beta}^{\rm HO}_{q}(\vec{r})=$ $\phi_{q}(\vec{r}),\qquad q\leq v=N-1,$ (61a)
${\beta}^{\rm HO}_{q}(\vec{r})=$ $\phi_{q}(\vec{r}),\qquad q\leq v-1=N-2.$ (61b)
This choice is supported by our intention to express the wave function $u(E;\vec{r})$ as an infinite series of oscillator functions. However, in many-body applications it will be difficult to select from a huge number of Slater determinants comprising the NCSM basis a reasonable set of a relatively small number of antisymmetrized SRFs with a given $J$ value. One, of course, can use some assumptions, e. g., the ideas of the RGM and define SRFs as a set of few relative motion functions times an antisymmetrized product of cluster wave functions supposing that the structure of colliding nuclei is unchanged during the collision. This is, of course, an approximation which works well in some cases but would be better to avoid for more general applications including the NCSM.
With the harmonic oscillator SRFs (61), we expect the expansion coefficients of the wave function (51) to be
$d_{n}(E)=\begin{cases}b_{n}(E)+a^{as}_{n}(E),&n=0,...\,,N-2,\\ a^{as}_{n}(E),&n=N-1,...\,,\infty,\end{cases}$ (62)
with $a^{as}_{n}(E)$ given by Eq. (53). The results of the ${}^{1}S_{0}$ phase shift calculations using the oscillator basis with ${\hbar\Omega=30}$ MeV truncated at $N_{{\max}}=16$ are presented in Fig. 2 in comparison with the exact phase shifts for this $NN$ interaction.
Figure 2: ${}^{1}S_{0}$ phase shifts obtained for various $N$ with $N_{\textrm{max}}=16$ and $\hbar\Omega=30$ MeV using the harmonic oscillator SRF. The inset in this and in figures below presenting ${}^{1}S_{0}$ phase shifts shows their low-energy behavior in a larger scale.
In the case of $N=\mathcal{\color[rgb]{0,0,0}N}=9$, the set of SRFs (61) results in the unphysical phase shifts of the same type as the set of eigenfunction SRFs (54) presented in Fig. 1. This is not surprising since both the functions (61a) and (54a) form a complete basis in the Hilbert space spanned by the oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$. Hence, any of the functions $\bar{\beta}^{\rm HO}_{q}(\vec{r})$ can be expressed as a superposition of the functions $\bar{\beta}^{\rm Efn}_{q}(\vec{r})$ and the set of equations (59) guarantees at any energy $E\neq E_{q}$ that
$\langle{\bar{\beta}_{q}^{\textrm{HO}}}|{u}(E;\vec{r})\rangle=0,\quad q=0,...\,,N-1$ (63)
with no coupling to the continuum and results in obtaining expansion coefficients $d_{n}(E)=0$ for all $n=0,$ , $\mathcal{\color[rgb]{0,0,0}N}-1$.
The situation improves dramatically when $N<\mathcal{\color[rgb]{0,0,0}N}$. The phase shifts $\delta_{0}(E)$ demonstrate clear convergence patterns with increasing $N$ and a good description of exact phase shifts is achieved even at relatively small values of $N=5$.
III.2 Proper application by coupling to the continuum
The success of the phase shift calculations with harmonic oscillator SRFs with $N<\mathcal{\color[rgb]{0,0,0}N}$ seems surprising. The problem is that we were using an inconsistent formulation of the problem which disappears in the case of $N<\mathcal{\color[rgb]{0,0,0}N}$ for the oscillator SRFs. In particular, our Hamiltonian (50) was confined to the Hilbert space spanned by the oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$. At the same time, we supposed that our wave function $u(E;\vec{r})$ has an infinite expansion in oscillator functions (51) where, according to Eq. (52), the expansion coefficients $d_{n}(E)=a^{as}_{n}(E)$ at $n\geq\mathcal{\color[rgb]{0,0,0}N}$. The asymptotic expansion coefficients $a^{as}_{n}(E)$ [see Eq. (53)] describing the free motion at energy $E$ and within the HORSE formalism fit the free Schrödinger equation BangHORSE ; Zaitsev1998
$\sum_{n^{\prime}=0}^{\infty}(T^{l}_{nn^{\prime}}-E\delta_{nn^{\prime}\!})\,a^{as}_{n^{\prime}}(E)=0,\qquad n=1,...\,,\,\infty,$ (64)
which, due to the tridiagonal structure of the kinetic energy matrix elements, Eqs. (41), reduces to the three-term recurrent relation
$T^{l}_{n,n-1}a^{as}_{n-1}(E)+(T^{l}_{nn}-E)a^{as}_{n}(E)+T^{l}_{n,n+1}a^{as}_{n+1}(E)=0.$ (65)
Therefore, according to our proposition regarding the structure of the wave function expansion coefficients, Eqs. (52)–(53), the Hamiltonian (50) should be extended to infinity in the oscillator space by the kinetic energy matrix elements,
$H=\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle V_{nn^{\prime}}\langle\phi_{n^{\prime}}|+\sum_{n,n^{\prime}=0}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|\\ \equiv\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle H^{tr}_{nn^{\prime}}\langle\phi_{n^{\prime}}|+\sum_{n,n^{\prime}=\mathcal{\color[rgb]{0,0,0}N}}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|\\ +|\phi_{\mathcal{\color[rgb]{0,0,0}N}-1}\rangle T^{l}_{\mathcal{\color[rgb]{0,0,0}N}-1,\mathcal{\color[rgb]{0,0,0}N}}\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}}|+|\phi_{\mathcal{\color[rgb]{0,0,0}N}}\rangle T^{l}_{\mathcal{\color[rgb]{0,0,0}N},\mathcal{\color[rgb]{0,0,0}N}-1}\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}-1}|,$ (66)
where the truncated Hamiltonian $H^{tr}$ given by Eq. (50) has the structure conventionally used in variational calculations like NCSM and its matrix elements $H^{tr}_{nn^{\prime}}$ include both potential energy matrix elements $V_{nn^{\prime}}$ and kinetic energy matrix elements $T^{l}_{nn^{\prime}}$ with $n,n^{\prime}=0$, … , $\mathcal{\color[rgb]{0,0,0}N}-1$, the term $\sum_{n,n^{\prime}=\mathcal{\color[rgb]{0,0,0}N}}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|$ describes the free motion (continuum) of colliding particles, and the last two terms in Eq. (66) describe the coupling between the continuum and the interaction region in the oscillator space.
The role of these interaction region–continuum coupling terms is very important and cannot be neglected. However, in the case of eigenfunction SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ and $\bar{\beta}_{q}^{\textrm{Efn}}(\vec{r})$, these coupling terms were completely suppressed since all matrix elements $\langle\bar{\beta}_{q}^{\textrm{Efn}}|H^{tr}|\phi_{\mathcal{\color[rgb]{0,0,0}N}}\rangle=\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}}|H^{tr}|\beta_{q}^{\textrm{Efn}}\rangle=0$. As a result, we arrived at an unphysical Eq. (60) for the phase shifts. The case of a “complete” set of harmonic oscillator SRFs $\bar{\beta}^{\rm HO}_{q}(\vec{r})$, Eq. (61), with $N=\mathcal{\color[rgb]{0,0,0}N}$ is equivalent to that of the eigenfunction SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ as discussed in the previous subsection. In the expansion of the wave function (51), all coefficients $d_{n}(E)=0$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$ and we have an analog of the scattering by an impenetrable sphere in the coordinate space with the unphysical phase shifts of the same type.
However in the case of an “incomplete” set of harmonic oscillator SRFs $\bar{\beta}^{\rm HO}_{q}(\vec{r})$ with $N<\mathcal{\color[rgb]{0,0,0}N}$, by solving Eq. (33), we obtain a set of functions $g_{n}(E)$ with $n=0$, … , $N-2$ in Eq. (52) and get non-zero expansion coefficients ${d_{n}(E)=a^{as}_{n}(E)}$ with $n\geq N-1$ which mimic the role of the last three terms in Eq. (66) though these terms were not explicitly included in the Hamiltonian of Eq. (50).
So, for the correct application of the Efros formalism with oscillator basis, one should use in the single-channel case the Hamiltonian (66). Regarding the SRFs $\beta_{q}(\vec{r})$, one can either use a set of $N-1$ oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{\color[rgb]{0,0,0}N}-1$, eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ of the Hamiltonian $H^{tr}$, or a set of ${N-1}$ linearly-independent superpositions of these functions which may be non-orthogonal. As for the functions $\bar{\beta}_{q}(\vec{r})$, we suggest to use ${N-1}$ functions $\beta_{q}(\vec{r})$ plus an additional function $\bar{\beta}_{N-1}(\vec{r})$ that guarantees the coupling to the continuum. It is most convenient to set it equal to $\phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r})$ which is coupled to a single function $\phi_{\mathcal{\color[rgb]{0,0,0}N}-1}(\vec{r})$ in the interaction region by the analytic matrix element $T^{l}_{\mathcal{\color[rgb]{0,0,0}N}-1,\mathcal{\color[rgb]{0,0,0}N}}$. That is, we choose
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}(\vec{r}),&0\leq q\leq N-2,\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r}),&q=N-1.\end{cases}$ (67)
We note that in the case of $N=\mathcal{\color[rgb]{0,0,0}N}$ (and, generally speaking, with $N>\mathcal{\color[rgb]{0,0,0}N}$) the version of the Efros formalism under discussion is equivalent to the HORSE formalism BangHORSE ; Zaitsev1998 ; Yamani , though these formalisms utilize different equations for calculating phase shifts, wave functions, etc. The advantage of the Efros approach is that it can use a smaller number $N$ of SRFs $\bar{\beta}_{q}(\vec{r})$ than the dimensionality of the truncated Hamiltonian matrix $\mathcal{\color[rgb]{0,0,0}N}$. The HORSE formalism provides exact solutions for the Hamiltonian Eq. (66) and its multichannel extensions including their versions with smoothed interactions [see Eqs. (68) and (75) below]. The Hamiltonian (66) is an approximation since the potential energy is truncated in the oscillator space, however the accuracy of this approximation is improved by increasing the truncation boundary ${\cal N}$. The Efros formalism with $N<{\cal N}~{}$ suggests an approximation of the HORSE and generates solutions which converge to the HORSE solutions with increasing $N$. We shall see below that one can obtain well-converged results with $N<\mathcal{\color[rgb]{0,0,0}N}$ and even with $N\ll\mathcal{\color[rgb]{0,0,0}N}$ in some cases.
Figure 3: ${}^{1}S_{0}$ phase shifts obtained by the HORSE method for various $N_{\textrm{max}}$ with $\hbar\Omega=30$ MeV.
Figure 4: ${}^{1}S_{0}$ phase shifts obtained by the HORSE method for various $\hbar\Omega$ at $N_{\textrm{max}}=24$.
III.3 Convergence of phase shifts and amplitudes
We start with examining the ${}^{1}S_{0}$ phase shifts within the Efros method with the SRF basis set with ${N=\mathcal{\color[rgb]{0,0,0}N}+1}$ which is equivalent to the HORSE method utilizing the same Hamiltonian (66), and investigate the phase shift convergence with $N_{\textrm{max}}$ at various $\hbar\Omega$. Figure 3 shows the convergence with $N_{\textrm{max}}$ at $\hbar\Omega=30$ MeV. It is seen that at $N_{\textrm{max}}=18$, we obtain a very good description of the phase shifts while at $N_{\textrm{max}}=24$, a complete convergence is achieved in a wide range of energies at this particular $\hbar\Omega$ value. Figure 4 shows the change of the phase shifts with $\hbar\Omega$ at $N_{\textrm{max}}=24$. We see that small $\hbar\Omega$ values are less optimal for convergence and henceforth, we use $\hbar\Omega=30$ MeV while studying the $np$ scattering in what follows.
Figure 5: ${}^{1}S_{0}$ phase shifts obtained by the Efros method using a) $N=4$ and $N_{\textrm{max}}=10$, b) $N=4$ and ${N_{\textrm{max}}=14}$, and c) $N=5$ and ${N_{\textrm{max}}=14}$ with both the harmonic oscillator and eigenfunction SRFs, ${\hbar\Omega=30}$ MeV.
Figure 6: ${}^{1}S_{0}$ phase shifts obtained by the Efros method with various $N$ at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV using the eigenfunction SRFs.
Figure 5 compares the ${}^{1}S_{0}$ phase shifts $\delta_{0}(E)$ obtained via the eigenfunction and harmonic oscillator SRFs for three different truncations, namely $N=4$, ${N_{\textrm{max}}=10}$, $N=4$, ${N_{\textrm{max}}=14}$, and $N=5$, $N_{\textrm{max}}=14$. We observe complicated convergence patterns. On the one hand, it is seen that switching from $N_{\max}=10$ to $N_{\max}=14$ at fixed $N=4$ improves the results with the harmonic oscillator SRFs but the results with the eigenfunction SRFs become even worse. On the other hand, switching from $N=4$ to $N=5$ at fixed $N_{\max}=14$ improves essentially the results with the eigenfunction SRFs but the results with the harmonic oscillator SRFs become slightly worse. However, generally, the results improve with increasing $N$ by more than one unit as well as with essential increase of $N_{\max}$.
In general, it is not obvious whether the harmonic oscillator SRFs or the eigenfunction SRFs perform better for the phase shifts. The lowest-energy results for the eigenfunction SRFs look better at the $N=4$ and $N_{\textrm{max}}=10$ truncations. However, the harmonic oscillator SRFs perform better at other energy intervals such as from $E=5$ MeV to $E=20$ MeV with that same truncation. The eigenfunction SRFs provide better convergence around ${E=45{-}50}$ MeV. Both approaches are comparable in the case of the ${N=5}$, $N_{\textrm{max}}=14$ truncation, though the harmonic oscillator SRFs perform better at the lowest energies up to approximately $E\approx 4$ MeV. From $E=4$ MeV to ${E=15}$ MeV, the eigenfunction SRFs result in a better description of the phase shifts, and the harmonic oscillator SRFs provide slightly better-converged phase shifts from $E=15$ MeV to ${E=35}$ MeV. From $E=35$ MeV to ${E=50}$ MeV, the results are virtually indistinguishable. The two sets of SRFs are comparable in terms of convergence, but in what follows we shall concentrate mostly on the case of the eigenfunction SRFs since they are more viable for many-body applications and suggest a more natural connection to the NCSM.
In Fig. 6 we investigate in more detail the convergence of $\delta_{0}(E)$ obtained by the Efros method with $N$ for $N_{\textrm{max}}=14$. We have noted a very reasonable convergence at $N=5$. However, at larger $N$, namely at $N=8$, the results start deviating significantly from the exact result. The irregularity of convergence is a consequence of the short-range behavior of the $NN$ interaction and a sudden drop-off of its matrix elements $V_{nn^{\prime}}$ at $n,n^{\prime}=\cal N$ due to the sharp truncation at relatively small $N_{\textrm{max}}$.
Figure 7: The effect of smoothing with various parameters $a$ on calculations of the ${}^{1}S_{0}$ phase shifts by the HORSE method with $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV.
Figure 8: ${}^{1}S_{0}$ phase shifts obtained by the HORSE method with various $N_{\textrm{max}}$ at $\hbar\Omega=30$ MeV with smoothing with parameter $a=2.5$.
To address the irregular convergence issue, we smooth the potential truncation as suggested in Refs. HungOne ; HungTwo by replacing in Eq. (66) the potential energy operator $\mathop{\sum}\limits_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle V_{nn^{\prime}}\langle\phi_{n^{\prime}}|$ by
$\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle\tilde{V}_{nn^{\prime}}\langle\phi_{n^{\prime}}|=\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle\sigma_{\mathcal{\color[rgb]{0,0,0}N}-1}^{n}V_{nn^{\prime}}\sigma_{\mathcal{\color[rgb]{0,0,0}N}-1}^{n^{\prime}}\langle\phi_{n^{\prime}}|,$ (68)
where
$\sigma_{\mathcal{\color[rgb]{0,0,0}N}-1}^{n}=\frac{1-e^{-\left(\!a\frac{n-\mathcal{\color[rgb]{0,0,0}N}}{\mathcal{\color[rgb]{0,0,0}N}}\!\right)^{2}}}{1-e^{-a^{2}}}$ (69)
and $a$ is a dimensionless parameter. The recommended values for the parameter $a$ are from the interval ${a=(2.5,10)}$. Figure 7 demonstrates the effect of smoothing on the phase shift calculations within the HORSE method with $N_{\textrm{max}}=14$. The parameter value of $a=2.5$ seems to be sufficient for our purposes and is henceforth used in our studies of $NN$ scattering.
Figure 9: ${}^{1}S_{0}$ phase shifts obtained by the Efros method with various $N$ at $N_{\textrm{max}}=14$ and $\hbar\Omega=30$ MeV using the eigenfunction SRFs and smoothing with parameter $a=2.5$.
The benefits of smoothing become more apparent in Figs. 8 and 9. We see that smoothing results in excellent convergence of the phase shifts within the HORSE method at $N_{\textrm{max}}=14$ as opposed to the $N_{\textrm{max}}=24$ needed for convergence without smoothing (see Fig. 3). In the case of the Efros method (Fig. 9), we obtain reasonable ${}^{1}S_{0}$ phase shifts with $N=5$ at $N_{{\max}}=14$ while $N=6$ provides very accurate results for $\delta_{0}(E)$ to at least $E=50$ MeV. Contrary to the case of non-smoothed interaction, a further increase of $N$ does not destroy the convergence.
It is important to check whether the proposed approach provides the correct scattering wave function, i. e., we reproduce the coefficients $d_{n}(E)$ in the wave function expansion, Eq. (51). We note that we can expect to obtain correct $d_{n}(E)$ values only in the case when we reproduce, at least with a reasonable accuracy, the scattering phase shifts $\delta_{l}(E)$ which enter Eq. (53) for the expansion coefficients $a^{as}_{n}(E)$. We present in Fig. 10 the convergence with $N$ of the coefficients $d_{n}(E)$ at the energy ${E=0.5}$ MeV where the ${}^{1}S_{0}$ phase shift obtained with smoothing at $N_{\max}=14$ and $\hbar\Omega=30$ MeV is close to the exact. The oscillator amplitudes $d_{n}(E)$ are seen to converge reasonably well. We note that the smoothing results in insignificant changes of the amplitudes $d_{n}(E)$ of the order of the line thickness if the phase shifts obtained with smoothing are very close to those obtained without the smoothing, but the amplitudes $d_{n}(E)$, of course, differ in the case when the smoothing results in the changes of the phase shift values.
Figure 10: Coefficients $d_{n}(E)$ of the wave function expansion (51) at energy ${E=0.5}$ MeV in the ${}^{1}S_{0}$ partial wave obtained with various $N$ at ${N_{\textrm{max}}=14}$ and $\hbar\Omega=30$ MeV with smoothing parameter $a=2.5$ using the eigenfunction SRFs.
In some cases, it may be useful to consider a basis that mixes the harmonic oscillator SRFs and the eigenfunction SRFs. For example, in many-body applications, it may happen that only a few lowest eigenstates are in the energy range of interest and those higher in energy do not significantly contribute to the low-energy phase shifts. In this case, it seems reasonable to introduce a few additional harmonic oscillator SRFs describing the relative motion of colliding fragments in the channel of interest. Therefore, we consider also a couple of sets of non-orthonormal linearly-independent SRFs combining both the harmonic oscillator and the eigenfunctions,
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...\,,q_{0},\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}-q+q_{0}}(\vec{r}),&q=q_{0}+1,...\,,N-2,\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r}),&q=N-1,\end{cases}$ (70)
and
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...\,,q_{0},\\ \phi_{q-q_{0}-1}(\vec{r}),&q=q_{0}+1,...\,,N-2,\\ \phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r}),&q=N-1.\end{cases}$ (71)
Here, $q_{0}+1$ is the number of low-lying eigenfunctions used. We note that for the case $N=\mathcal{\color[rgb]{0,0,0}N}+1$, the HORSE method BangHORSE results are reproduced regardless of the basis combination used. For both Eqs. (70) and (71), the set of SRFs $\{\beta_{q^{\prime}}\},$ $q^{\prime}=0,...\,,N-2$ appearing explicitly in Eq. (19) is the same as the set of SRFs $\{\bar{\beta}_{q}\}$ for $q=0,...\,,N-2$ but it excludes $\bar{\beta}_{N-1}=\phi_{\mathcal{\color[rgb]{0,0,0}N}}(\vec{r})$.
In Fig. 12 we present the ${}^{1}S_{0}$ phase shifts obtained with the SRF set of Eq. (70) with $N_{\textrm{max}}=14$, $\hbar\Omega=30$ MeV, and $N=6$. Vertical lines correspond to the energies of eigenstates of the Hamiltonian $H^{tr}$ at which the phase shifts are well-described AMShirokovHORSE2016 . We see that at least $q_{0}=3$ is needed to get reasonable phase shifts in a wide range of energies while with $q_{0}=4$ the results for the phase shifts are excellent.
Figure 11: ${}^{1}S_{0}$ phase shifts obtained by the Efros method with various $q_{0}$ using the hybrid SRF set given by Eq. (70), $N_{\textrm{max}}=14$, $N=6$, $\hbar\Omega=30$ MeV, and the smoothing parameter $a=2.5$. Vertical lines show the energies of eigenstates of the Hamiltonian $H^{tr}$.
Figure 12: The same as Fig. 12 but using the hybrid SRF set given by Eq. (71) and $N_{\max}=12$.
Figure 12 presents the same results but obtained with the SRF set of Eq. (71) and smaller ${N_{\textrm{max}}=12}$. We see that in this case, we get reasonable results even with $q_{0}=0$ and much more accurate with $q_{0}=1$. Thus it is preferable to use in applications the SRF set of Eq. (71).
IV Two-Channel Scattering
We have shown that the Efros method is consistent and leads to a convergence of single-channel scattering calculations using a limited basis space obtained by diagonalizing the truncated Hamiltonian $H^{tr}$. We now extend our application to a system with two open channels. In particular, we investigate $np$ scattering in the ${}^{3}SD_{1}$ channel using the Daejeon16 $NN$ potential and eigenfunction SRFs.
IV.1 Hamiltonian, $S$ matrix and Stapp parameterization for $np$ scattering in coupled partial waves
The Hamiltonian of the $np$ system in the coupled partial waves can be written as
$H={\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle V_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|+{\color[rgb]{0,0,0}\sum_{i=1,2}}\>\sum_{n,n^{\prime}=0}^{\infty}|\phi_{ni}\rangle T^{l_{i}}_{nn^{\prime}}\langle\phi_{n^{\prime}i}|\equiv{\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle H^{tr}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|\\ +\sum_{i=1,2}\>\sum_{n,n^{\prime}=\mathcal{\color[rgb]{0,0,0}N}_{i}}^{\infty}|\phi_{ni\!}\rangle T^{l_{i}}_{nn^{\prime}}\langle\phi_{n^{\prime}i}\!|\!+\sum_{i=1,2}\!\big{[}|\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1,i\!}\rangle T^{l_{i}}_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1,\mathcal{\color[rgb]{0,0,0}N}_{i}}\!\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}i}\!|+|\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}i}\rangle T^{l_{i}}_{\mathcal{\color[rgb]{0,0,0}N}_{i},\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\langle\phi_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1,i}|\big{]}\!.$ (72)
We define the truncation of the potential energy matrix by the maximal oscillator quanta $N_{\max}$, i. e., ${2n+l_{i}}\leq N_{\max}$ and $\mathcal{\color[rgb]{0,0,0}N}_{i}=\frac{1}{2}(N_{\max}-l_{i})+1$, $i=1,2$.
The smoothing of the interaction can be also applied in the two-channel case by introducing the potential energy operator
$\tilde{V}\equiv\sum_{i,j=1,2}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle\tilde{V}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|=\sum_{i,j=1,2}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle\sigma_{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}^{n}V_{ni,n^{\prime}j}\sigma_{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}^{n^{\prime}}\langle\phi_{n^{\prime}j}|.$ (73)
Smoothing also improves the convergence in the two-channel case. We are not presenting results prior to smoothing but only present below results of calculations with the smoothed interactions. That means that we use the truncated Hamiltonian
$H^{tr}\equiv\tilde{H}^{tr}={\color[rgb]{0,0,0}\sum_{i,j=1,2}}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle H^{tr}_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|$ (74)
with matrix elements
$H^{tr}_{ni,n^{\prime}j}\equiv\tilde{H}^{tr}_{ni,n^{\prime}j}=\langle\phi_{ni}|T+\tilde{V}|\phi_{n^{\prime}j}\rangle,\quad n<{\cal N}_{i},\ n^{\prime}<{\cal N}_{j}.$ (75)
For the two-channel problem, it is convenient to solve Eq. (22) in the $S$-matrix representation. We define the set of SRFs $\{\beta_{q}\}$ as a set of $N-2$ eigenfunctions of $H^{tr}\!$,
${\beta}_{q}(\vec{r})=\beta_{q}^{\textrm{Efn}}(\vec{r}),\quad q=0,...,N-3.$ (76)
Note, $N=v+2$ is the number of unknowns in each channel and the total number of independent unknowns is ${{\mathbb{\color[rgb]{0,0,0}N}}=2N-1\color[rgb]{0,0,0}=2v+3}$. We first solve the full set of equations for the $i=1$ entrance channel with the set $\{\bar{\beta}_{q}^{1}\}$ constituting a set of $N-2$ eigenfunctions of $H^{tr}$ plus two additional oscillator functions that are used exclusively for orthogonalization and guarantee coupling to the continuum but do not appear explicitly in Eq. (1):
${\color[rgb]{0,0,0}\bar{\beta}_{q}^{1}(\vec{r})}=\begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...,N-3,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{1}1}(\vec{r}),&q=N-2,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{2}2}(\vec{r}),&q=N-1.\end{cases}$ (77)
Here,
$\beta_{q}^{\textrm{Efn}}(\vec{r})=\sum_{q^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{1}-1}a_{qq^{\prime}}^{l_{1}}\,{\color[rgb]{0,0,0}\phi_{q^{\prime}1}(\vec{r})}+\sum_{q^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{2}-1}a_{qq^{\prime}}^{l_{2}}\,\color[rgb]{0,0,0}\phi_{q^{\prime}2}(\vec{r}).$ (78)
Next, we solve the reduced set of equations for the ${i=2}$ entrance channel. Since there is one less unknown, namely, the off-diagonal $S$-matrix element ${\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)}$, we use for the set $\{\bar{\beta}_{q}^{2}\}$ a subset of $\{\bar{\beta}_{q}^{1}\}$ that includes one less function. The choice of the function to be omitted for the orthogonalization in the $i=2$ channel is arbitrary. We can omit, e. g., $\beta_{N-3}^{\textrm{Efn}}(\vec{r})$, and define the set $\{\bar{\beta}_{q}^{2}\}$ as
${\color[rgb]{0,0,0}\bar{\beta}_{q}^{2}(\vec{r})}=\begin{cases}\beta_{q}^{\textrm{Efn}}(\vec{r}),&q=0,...,N-4,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{1}1}(\vec{r}),&q=N-3,\\ \color[rgb]{0,0,0}\phi_{\mathcal{\color[rgb]{0,0,0}N}_{2}2}(\vec{r}),&q=N-2.\\ \end{cases}$ (79)
We note that the set of SRFs $\{\beta_{q}\}$ appearing explicitly in Eq. (1) and defined by Eq. (76) is identical in both channels.
Since we have specified the SRFs, we can write down the coefficients $d_{n,ij}(E)$ of the continuum wave function $u_{i}(E;\vec{r})$ expansion in oscillator functions [see Eq. (12)]. For the $i=1$ entrance channel, the wave function $u_{1}(E;\vec{r})$ has a “strong” component $\color[rgb]{0,0,0}\sum_{n=0}^{\infty}d_{n,11}(E)\,\phi_{n1}(\vec{r})$ in the channel 1 and a “weak” component $\color[rgb]{0,0,0}\sum_{n=0}^{\infty}d_{n,12}(E)\,\phi_{n2}(\vec{r})$ in the channel 2 obtained from the set of $i=1$ linear equations. The respective expansion coefficients are
$d_{n,11}(E)\!=\!\begin{cases}a_{n,11}^{as}(E)+g_{n,11}(E),&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}_{1}-1,\\ a_{n,11}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{1},...\,,\infty\end{cases}$ (80)
and
$d_{n,12}(E)\!=\!\begin{cases}a_{n,12}^{as}(E)+g_{n,12}(E),\!\!&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}_{2\!}-1,\\ a_{n,12}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{2},...\,,\infty.\end{cases}$ (81)
For the $i=2$ entrance channel, the wave function $u_{2}(E;\vec{r})$ has a “weak” component $\color[rgb]{0,0,0}\sum_{n=0}^{\infty}d_{n,21}(E)\,\phi_{n1}(\vec{r})$ in the channel 1 and a “strong” component $\color[rgb]{0,0,0}\sum_{n=0}^{\infty}d_{n,22}(E)\,\phi_{n2}(\vec{r})$ in the channel 2 with the expansion coefficients
$d_{n,21}(E)\!=\!\begin{cases}a_{n,21}^{as}(E)+g_{n,21}(E),\!\!&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}_{1}-1,\\ a_{n,21}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{1},...\,,\infty\end{cases}$ (82)
and
$d_{n,22}(E)=\begin{cases}a_{n,22}^{as}(E)+g_{n,22}(E),&n=0,...,\mathcal{\color[rgb]{0,0,0}N}_{2}-1,\\ a_{n,22}^{as}(E),&n=\mathcal{\color[rgb]{0,0,0}N}_{2},...,\infty.\end{cases}$ (83)
In Eqs. (80)–(83),
$\!\!a_{n,ij}^{as}(E)$ $=C^{-}_{nl_{i}}(k)\,\delta_{ij}-\mathbb{S}_{ji}(E)\,C^{+}_{nl_{j}}(k),\!\!$ $i,j=1,2$ (84)
$\!\!g_{n,ij}(E)$ $=\sum_{n^{\prime}=0}^{N-3}b_{n^{\prime}}^{i}(E)\,a_{n^{\prime}n}^{l_{j}},\!$ $i,j=1,2.$ (85)
In the case of the $np$ scattering, it is conventional to express the $S$ matrix in the Stapp parameterization Stapp where the phase shifts $\delta_{l_{i}}=\delta_{l_{i}}(E)$ are introduced in each of the partial waves and the mixing parameter $\epsilon=\epsilon(E)$. These parameters are related to the $S$ matrix as
$\mathbb{S}(E)=\begin{pmatrix}e^{2i\delta_{l_{1}}}\textrm{cos}(2\epsilon)&ie^{i(\delta_{l_{1}}+\delta_{l_{2}})}\textrm{sin}(2\epsilon)\\[3.00003pt] ie^{i(\delta_{l_{1}}+\delta_{l_{2}})}\textrm{sin}(2\epsilon)&e^{2i\delta_{l_{2}}}\textrm{cos}(2\epsilon)\end{pmatrix}\!.$ (86)
IV.2 Two-channel results
Figure 13: a) ${}^{3}S_{1}$ phase shifts $\delta_{0}$, b) ${}^{3}D_{1}$ phase shifts $\delta_{2}$, c) ${}^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the HORSE method with $\hbar\Omega=30$ MeV and various $N_{\textrm{max}}$; the smoothing parameter $a=2.5$. The inset in (c) displays $\epsilon$ in a larger scale on the y-axis.
We present in Fig. 13 the results of calculations of the phase shifts and mixing parameter $\epsilon$ in the ${{}^{3}SD}_{1}$ coupled partial waves of $np$ scattering using the HORSE method with $\hbar\Omega=30$ MeV and smoothing parameter $a=2.5$. The ${}^{3}S_{1}$ phase shifts $\delta_{0}$ convergence is fast and at $N_{\max}=6$ we have accurate results. The ${}^{3}D_{1}$ phase shift $\delta_{2}$ convergence is also good and starting from $N_{\max}=6$ the phase shifts are reasonable (note the expanded scale in Fig. 13b as compared with the scale in Fig. 13a) while with ${N_{\max}=10}$ and $N_{\max}=12$, we obtain an excellent description of the phase shifts. The mixing parameter $\epsilon$ is also reasonably well-described at $N_{\max}=12$ (note, the scale in Fig. 13c is further expanded), we obtain small deviations of $\epsilon$ from the exact values at $N_{\max}=16$ and a very accurate description of the mixing parameter at $N_{\max}=20$.
If we use a “complete” set of SRFs $\{\beta_{q}\}$ in Eq. (1) which is enough to form a complete basis in the subspace of the oscillator functions $\phi_{ni}(\vec{r})$ with $n=0,1,...\,,{\cal N}_{i}$, $i=1,2$, the Efros approach is equivalent to the HORSE method. This set of SRFs $\{\beta_{q}\}$ may be arbitrary: it is only important that the functions $\beta_{q}$ are linearly-independent. In this case, one can use the over-complete Eq. (27) which suggest additional relations for independent calculations of the unknowns that should be equal such as $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$ in the case of two-channel scattering. As an example we consider the $S$-matrix calculations at the energy of $E=0.5$ MeV in the coupled ${}^{3}SD_{1}$ partial waves of $np$ scattering with ${\hbar\Omega=30}$ MeV and smoothing parameter $a=2.5$. We employ the set of $H^{tr}$ eigenfunction SRFs for $N_{\textrm{max}}=12$ and $N=15$, i. e., all of the 13 $H^{tr}$ eigenfunctions are used in the set of the SRFs $\{\beta_{q}\}$ and two additional oscillator functions used exclusively for orthogonalization are included in the set $\{\bar{\beta}_{q}\}$ as in Eq. (77). We obtain the $S$ matrix
$\mathbb{S}(E)=\begin{pmatrix}0.4193-0.9078i&-0.0022-0.0035i\\[3.00003pt] -0.0022-0.0035i&1-0.0002i\\[3.00003pt] \end{pmatrix}\!,$
which is unitary and $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)$. Since we can switch from the “complete” set of eigenfunction SRFs to an equivalent set of harmonic oscillator SRFs via a unitary transformation, both sets of SRFs yield the same solutions. Thus the “complete” set of SRFs $\{\beta_{q}\}$ guarantees the consistency of relations between unknowns following from Eq. (27).
As it was already noted, in the above case of the “complete” set of SRFs $\{\beta_{q}\}$, the Efros and HORSE methods are equivalent. Since the “complete” set of SRFs in current ab initio nuclear structure applications is extremely large, it is more interesting to investigate the essential simplification suggested within the Efros method as compared to the HORSE method which makes it possible to use a reduced set of SRFs. Let us consider the same case as above, but for $N=12$. In this case, Eq. (77) results in the $S$ matrix
$\mathbb{S}(E)=\begin{pmatrix}0.4169-0.909i&-0.0021-0.0033i\\[3.00003pt] -0.0016-0.0025i&1-0.0002i\\[3.00003pt] \end{pmatrix}$
with broken unitarity and $\mathbb{S}_{12}(E)\neq\mathbb{S}_{21}(E)$. In the case of a reduced set of SRFs $\{\beta_{q}\}$, one is pushed to use Eq. (33) and impose the $S$-matrix symmetry either by calculating $\mathbb{S}_{21}(E)$ and setting $\mathbb{S}_{12}(E)$ equal to it or vice versa.
Figure 14: a) ${}^{3}S_{1}$ phase shifts $\delta_{0}$, b) ${}^{3}D_{1}$ phase shifts $\delta_{2}$, c) ${}^{3}SD_{1}$ mixing parameter $\epsilon$ obtained by the Efros method with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and with smoothing parameter ${a=2.5}$ using the eigenfunction SRFs and ${}^{3}S_{1}$ as the first channel with $i=1$, i. e., imposing the symmetry by setting $\mathbb{S}_{12}(E)$ equal to $\mathbb{S}_{21}(E)$. See text for a detailed explanation of the label for each curve.
Figure 15: The same as Fig. 14 but using the ${}^{3}D_{1}$ as the first channel with $i=1$.
In the case of a limited set of SRFs, if we use the ${}^{3}S_{1}$ as the first channel corresponding to $i=1$ and set $l_{1}=0$ and $l_{2}=2$ [i. e., first solving the complete set of the ${}^{3}S_{1}$ channel equations, obtaining $\mathbb{S}_{21}(E)$, and next solving the reduced set of the ${}^{3}D_{1}$ equations imposing the symmetry by setting $\mathbb{S}_{12}(E)$ equal to $\mathbb{S}_{21}(E)$], we get results different from those obtained using the ${}^{3}D_{1}$ as the first channel corresponding to $i=1$ and setting $l_{1}=2$ and $l_{2}=0$ (i. e., solving first the complete set of equations in the ${}^{3}D_{1}$ partial wave and next switching to the reduced set of equations in the ${}^{3}S_{1}$ partial wave). We compare these approaches in Figs. 14 and 15 where we present, respectively, the results using the ${}^{3}S_{1}$ and the ${}^{3}D_{1}$ channel as the first.
The results for the phase shifts and the mixing parameter $\epsilon$ presented in Figs. 14 and 15 were obtained using $H^{tr}$ eigenfunctions as SRFs, $\hbar\Omega=30$ MeV, and truncation of the Hamiltonian $H^{tr}$ at $N_{\max}=12$ with the smoothing parameter $a=2.5$. We prefer to use the lowest-lying eigenfunctions of $H^{tr}$ in calculations. Let us introduce the notation used in the label for Fig. 14 (which is the same as for Fig. 15) by way of an explicit example. The $N=5\,\{0,1,2\}$ in the figure legends means that we use 5 unknowns in the $i=1$ partial wave [3 SRF coefficients $b^{2}_{q}(E)$, $q=0,1,2$ plus 2 $S$-matrix elements $\mathbb{S}_{11}(E)$ and $\mathbb{S}_{21}(E)$], 4 unknowns in the $i=2$ partial wave [3 SRF coefficients $b^{1}_{q}(E)$, $q=0,1,2$ plus 1 $S$-matrix element $\mathbb{S}_{22}(E)$], that is a total of $\color[rgb]{0,0,0}\mathbb{N}=9$ independent unknowns, and take as the SRFs the lowest (ground state) eigenfunction $\beta_{0}^{\textrm{Efn}}(\vec{r})$, $\beta_{1}^{\textrm{Efn}}(\vec{r})$, and $\beta_{2}^{\textrm{Efn}}(\vec{r})$. We omit ${\beta}_{2}^{\textrm{Efn}}(\vec{r})$ in the set $\{\bar{\beta}_{q}^{2}\}$ used for the orthonormalization of the reduced set of equations in the $i=2$ channel, but not in the set $\{\bar{\beta}_{q}^{1}\}$ used in the equations in the $i=1$ channel.
In the case of using the ${{}^{3}S}_{1}$ as the first channel presented in Fig. 14, the calculations with $N=5$ provide reasonable results for $\color[rgb]{0,0,0}\delta_{0}$ at low energies and for $\color[rgb]{0,0,0}\epsilon$ in a larger energy interval but a completely incorrect behavior of $\color[rgb]{0,0,0}\delta_{2}$. It is interesting to note that the results for $\delta_{0}$ at $N=5$ start deviating from the exact curve at energies around $E=11{-}12$ MeV. This roughly corresponds to the second excited state eigenenergy of $H^{tr}$ at $E_{2}=11.457$ MeV. The origin of strange unreasonable $\delta_{2}$ behavior is unclear. In the case of $N=6$, $\delta_{0}$ describes well the exact result while $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$ are reasonably accurate up to approximately $E=25$ MeV and $E=20$ MeV, respectively. In calculations with ${N=7}$ and $8$, we see emerging numerical problems. For example, a jump of $\delta_{0}$ around ${E=3}$ MeV occurs which is more pronounced in the $N=7$ than in the $N=8$ calculations. Except for these spikes, the $\delta_{0}$ convergence is reasonable. The calculations of $\delta_{2}$ with $N=7$ fail completely and with $N=8$ fail up to approximately 12 MeV but describe well the exact results at larger energies. The $\epsilon$ calculations with $N=7$ and $8$ provide close results which are reasonable at energies above 16 MeV but unreasonable at lower energies.
More reasonable results are obtained if setting the ${}^{3}D_{1}$ as the first channel (see Fig. 15). With the same parameters ($\hbar\Omega$, $N_{\max}$, etc.) we get good results with $N=5$ and $N=6$ at least up to $E=16$ MeV for all quantities. The results are still accurate in cases of $N=7$ and $N=8$, though $\color[rgb]{0,0,0}\delta_{0}$ obtained with $N=8$ shows some oscillation around $E\approx 12$ MeV and $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$ have jumps around $\color[rgb]{0,0,0}E\approx 3$ MeV in cases of both $N=7$ and $N=8$.
Figure 16: ${}^{3}SD_{1}$ channel: a) $|\textrm{det}(A)|$, b) $\textrm{Arg}(\textrm{det}(A))$ obtained in calculations with $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and the ${{}^{3}D}_{1}$ partial wave as the first channel with and without the basis rearrangement. The vertical line corresponds to the first excited state eigenenergy of $\color[rgb]{0,0,0}H^{tr}\!$.
Figure 17: a) ${}^{3}S_{1}$ phase shifts $\delta_{0}$, b) ${}^{3}D_{1}$ phase shifts $\delta_{2}$, c) ${}^{3}SD_{1}$ mixing parameter $\epsilon$ obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and with smoothing parameter ${a=2.5}$ using the eigenfunction SRFs and ${}^{3}S_{1}$ as the first channel with $i=1$ and rearranging the set of SRFs to eliminate the vanishing point in the determinant of $A$. This elimination is checked over a range of $E$ from 0 to 50 MeV.
Figure 18: The same as Fig. 17 but using ${}^{3}D_{1}$ as the first channel with $i=1$.
Figure 19: The same as Fig. 18 but calculations are performed with $N_{\max}=20$.
Figure 20: Absolute magnitudes of ${{}^{3}SD_{1}}$ wave function amplitudes at the energy of $E=0.5$ MeV obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and ${}^{3}D_{1}$ as the first channel with $i=1$: a) $|d_{n,22}(E)|$ (“strong” ${}^{3}S_{1}$ component), b) $|d_{n,21}(E)|$ (“weak” ${}^{3}S_{1}$ component), c) $|d_{n,12}(E)|$ (“weak” ${}^{3}D_{1}$ component), d) $|d_{n,11}(E)|$ (“strong” ${}^{3}D_{1}$ component).
Figure 21: Arguments of ${{}^{3}SD_{1}}$ wave function amplitudes at the energy of $E=0.5$ MeV obtained with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and ${}^{3}D_{1}$ as the first channel with $i=1$: a) $\textrm{Arg}(d_{n,22}(E))$ (“strong” ${}^{3}S_{1}$ component), b) $\textrm{Arg}(d_{n,21}(E))$ (“weak” ${}^{3}S_{1}$ component), c) $\textrm{Arg}(d_{n,12}(E))$ (“weak” ${}^{3}D_{1}$ component), d) $\textrm{Arg}(d_{n,11}(E))$ (“strong” ${}^{3}D_{1}$ component).
Figure 22: $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering supported by the Woods–Saxon potential: a) resonance energy $E_{r}$ and b) resonance width $\Gamma$ obtained by locating the $S$-matrix pole within the HORSE method using various $N_{\max}$ truncations, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$.
Figure 23: $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering supported by the Woods–Saxon potential: a) resonance energy $E_{r}$ and b) resonance width $\Gamma$ obtained by locating the $S$-matrix pole within the Efros method with eigenfunction SRFs using various $N$, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ in comparison with the exact values and with the HORSE method results.
Figure 24: Deuteron bound state energy $E_{b}$ obtained with Daejeon16 $NN$ interaction using various $N_{\textrm{max}}$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ by the HORSE method and by the direct diagonalization of $H^{tr}$.
Figure 25: Deuteron bound state energy $E_{b}$ obtained with Daejeon16 $NN$ interaction by locating the $S$-matrix pole via the Efros method using the eigenfunction SRFs, various $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$ in comparison with the exact value and with the HORSE method result.
Figure 26: Absolute magnitudes of expansion coefficients $|\breve{d}_{n,i}(E_{b})|$ of the deuteron wave function obtained by locating the $S$-matrix pole within the Efros method with different $N$ using eigenfunction SRFs, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$: a) ${}^{3}S_{1}$ components, b) ${}^{3}D_{1}$ components. The deuteron energy obtained with each $N$ is given in the legend in parentheses. We present also for comparison the exact values of $|\breve{d}_{n,i}(E_{b})|$ ($E_{b}=-2.225$ MeV) and those values obtained by the HORSE method. Symbols are often obscured by other symbols especially in panel (a).
Figure 27: a) Energy $E_{r}$ and b) width $\Gamma$ of the resonance in the two-channel Noro–Taylor model Norro1980 in HORSE calculations with $\Omega=1.5$ au, smoothing parameter $a=5$, and various $N_{\max}$.
Figure 28: a) Energy $E_{r}$ and b) width $\Gamma$ of the resonance in the two-channel Noro–Taylor model Norro1980 in calculations within Efros method with eigenfunction SRFs, $\Omega=1.5$ au, $N_{\max}=20$, smoothing parameter $a=5$, and various $N$ as compared with the exact results and with the HORSE method results.
The origin of these spikes is illustrated in Fig. 16, where $|\det(A)|$, the absolute magnitude of the matrix $A$ determinant, and $\textrm{Arg}\big{(}\textrm{det}(A)\big{)}$, the determinant argument, are plotted as functions of $E$ in the case of $N=7$ and using the ${}^{3}D_{1}$ as the first channel. If we use the set of the SRF eigenfunctions with increasing eigenvalues, $N=7\,\{0,1,2,3,4\}$, there is a clearly seen discontinuity of $\textrm{Arg}\big{(}\det(A)\big{)}$ around $E=3$ MeV which coincides with the spikes in Fig. 15 for $\color[rgb]{0,0,0}\delta_{2}$ and $\color[rgb]{0,0,0}\epsilon$. To fix this, we simply replace one of the SRFs $\beta_{q}^{\textrm{Efn}}(\vec{r})$ and change their ordering which results in replacing the omitted function $\bar{\beta}_{q}^{\textrm{Efn}}(\vec{r})$ in the orthonormalization of the $i=2$ equations. For example, the ordering $\{0,1,4,6,2\}$ fixes the discontinuities in $\textrm{Arg}(\textrm{det}(A))$ and the spikes that they cause.
This adjustment of the set of SRFs improves the results as demonstrated in Figs. 17 and 18 where modified sets of SRFs are indicated by the red colors in the legends. It is seen that with $N=8\,\{0,1,2,6,4,3\}$ in the case of ${{}^{3}S}_{1}$ used as the first channel or with $N=8\,\{0,1,2,3,4,6\}$ in the case of ${{}^{3}D}_{1}$ used as the first channel, we obtain an excellent description of both phase shifts $\delta_{0}$ and $\delta_{2}$ but an inaccurate description of the mixing parameter $\epsilon$. It is interesting that in the case of the first channel assigned to the ${{}^{3}S}_{1}$ partial wave, the best $\epsilon$ description is obtained with the smallest number of SRFs ${N=5\,\{0,2,1\}}$ when the phase shifts $\delta_{0}$ and $\delta_{2}$ are well-described in relatively small energy intervals, up to approximately 15 and 30 MeV, respectively (see Fig. 17). It is seen that the convergence trends with an increasing number of SRFs $N$ is complicated and non-monotonic in the two-channel case. We have not yet found a systematic way to change the SRF set ordering to eliminate the vanishing of the $A$ determinant. Generally, our preference is to retain the lowest-lying eigenfunctions since they are important in the formation of scattering observables at low energies.
As was already mentioned, the Efros method with a “complete” set of SRFs becomes equivalent to the HORSE approach. The description of the phase shifts with increasing $N$ approaches, though non-monotonically, to that of the HORSE method. As is seen in Fig. 13, the HORSE description of the scattering parameters, in particular, of the mixing parameter $\epsilon$, is still not ideal at $N_{\max}=12$ with $\hbar\Omega=30$ MeV. Therefore, to obtain a good description of all scattering observables of $np$ scattering in the ${{}^{3}SD}_{1}$ coupled partial waves in a wide energy range, one needs to perform calculations with a larger $N_{\max}$ value. We present in Fig. 19 the results of calculations with $N_{\max}=20$, with the rest of the parameters the same as in Fig. 18. It is seen that we get an excellent description of all scattering observables up to the energy of 50 MeV if we use a large enough set of SRFs $N=11\,\{0,1,2,3,4,5,6,7,8\}$.
As in the single-channel case, one cannot expect a good description of the wave function if the scattering phase shifts and mixing parameter are not reproduced. We present in Figs. 20 and 21 absolute magnitudes and arguments of the wave function expansion coefficients $d_{n,ij}(E)$ respectively, at the energy of $E=0.5$ MeV where we obtain $\delta_{0}$, $\delta_{2}$, and $\epsilon$ close to exact value in calculations with different $N$, $N_{\textrm{max}}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=2.5$ using the eigenfunction SRFs and ${}^{3}D_{1}$ as the first channel with $i=1$. We observe a fast convergence of the amplitudes even with small $N$ values.
In a wider range of energies, the wave function expansion coefficients $d_{n,ij}(E)$ obtained in calculations with $N_{\max}=12$ start deviating from the exact values because the phase shift $\delta_{0}$, $\delta_{2}$ and the mixing parameter $\epsilon$ descriptions are not accurate. However, the description of expansion coefficients $d_{n,ij}(E)$ may be improved by switching to larger $N_{\max}$ values. For example, in calculations with $N_{\max}=20$, we obtain excellent expansion coefficients $d_{n,ij}(E)$ up to an energy of 50 MeV using the set of SRFs $N=11\,\{0,1,2,3,4,5,6,7,8\}$, which provides an accurate description of scattering parameters as is seen in Fig. 19.
V ${S}$-matrix poles
We present and discuss calculations of the $S$-matrix poles within the HORSE and Efros approaches using three simple examples. In particular, we explore a resonance in $n\alpha$ scattering using the Woods–Saxon interaction with parameters fitted in Ref. WoodsSaxon1979 . Next, we consider an $S$-matrix pole associated with a bound state using as an example the deuteron bound state supported by the Daejeon16 $NN$ interaction. Finally, we analyze a narrow resonance in a model two-channel problem originally presented in Ref. Norro1980 and used as a benchmark for different numerical techniques in Refs. Moiseyev1990 ; Rakityansky2006 .
V.1 Single-channel resonance $S$-matrix pole
To describe the $n\alpha$ scattering with the Woods–Saxon interaction fitted in Ref. WoodsSaxon1979 , we use the reduced mass of $m=\frac{4}{5}m_{N}=751.136~{}{\textrm{MeV}}/{c^{2}}$, ${\hbar\Omega=30}$ MeV, and the smoothing parameter $a=5$. We investigate a resonance in the ${}^{2}P_{\frac{3}{2}}$ partial wave by locating numerically the corresponding $S$-matrix pole. Figure 22 shows the convergence with $N_{\textrm{max}}$ of the resonance energy $E_{r}$ and width $\Gamma$ obtained by the HORSE method which is equivalent to the Efros method in the case of a “complete” set of SRFs as discussed above; thus, to locate the $S$-matrix pole, one needs to solve Eq. (46) in the complex momentum plane. We see that at $N_{\textrm{max}}=10$, we obtain results very close to the exact values of $E_{r}=0.837$ MeV and $\Gamma=0.780$ MeV (see Refs. AMShirokovHORSE2016 ; BlokhintsevSSHorse ). At larger $N_{\max}$, there are small oscillations of both $E_{r}$ and $\Gamma$ around the exact values, and starting from $N_{\max}=20$, the energy and width are stable and coincide with the exact values up to at least three significant digits.
In the search for the $S$-matrix pole by the Efros method, which can use a smaller number of eigenstates, we perform the calculations with the set of the SRF eigenfunctions with increasing eigenvalues $N\,\{0,1,2,...\}$ at ${N_{\textrm{max}}=12}$ and compare with the result obtained within the HORSE method. The results are presented in Fig. 23. Both $E_{r}$ and $\Gamma$ exhibit oscillations, but note that the scale in Fig. 23 is significantly expanded as compared with that of Fig. 22. The energy and width are seen to converge with $N$ reasonably fast to the exact values and starting from $N=5$, differ from the exact values by less than 5%.
V.2 Bound state $S$-matrix pole
We next search for the $S$-matrix pole in the coupled ${}^{3}SD_{1}$ partial waves associated with the deuteron bound state energy $E_{b}$ using the Daejeon16 $NN$ interaction smoothed with the parameter $a=5$ and $\hbar\Omega=30$ MeV. We first utilize the HORSE method and present in Fig. 24 the convergence of $E_{b}$ with $N_{\textrm{max}}$. The deuteron ground state energy is seen to converge much faster to the exact value than the results of the direct diagonalization of the truncated Hamiltonian $H^{tr}$. In particular, in the case of the direct diagonalization, approximately $N_{\textrm{max}}=50$ is needed to achieve the accuracy of three significant digits. In contrast, by the $S$-matrix pole location via the HORSE method, we obtain the same accuracy at ${N_{\textrm{max}}=14}$, and the accuracy of four significant digits is achieved at $N_{\max}=20$.
We obtain also a fast convergence by locating the $S$-matrix pole using the Efros method which utilizes a limited number of SRF eigenfunctions with increasing eigenvalues $N\>\{0,1,2,...\}$. The results for the ground state energy $E_{b}$ obtained with $N_{\textrm{max}}=12$ are presented in Fig. 25. $E_{b}$ exhibits jumps and instability as $N$ increases up to $N=12$; note, however, an extended scale in Fig. 25. In particular, even at small $N$ starting from $N=3$, the fluctuations of $E_{b}$ never deviate by more than 4% from the exact value.
In the case of bound states, there are no entrance (exit) channels and the solution is unique and independent from the selection of a channel as the first channel. The expansion of the bound state wave function in the series of oscillator functions takes the form
$u(E_{b};\vec{r})=\sum_{i=1}^{w}\sum_{n=0}^{\infty}\breve{d}_{n,i}(E_{b})\,\phi_{n{i}}(\vec{r}).$ (87)
We present in Fig. 26 $\breve{d}_{n,i}(E_{b})$ of the deuteron wave function obtained by the Efros method with different $N$ using eigenfunction SRFs, $N_{\max}=12$, $\hbar\Omega=30$ MeV, and smoothing parameter $a=5$. We see a high accuracy of expansion coefficients calculated even with small $N$.
We note that with the diagonalization of the truncated Hamiltonian matrix $H^{tr}$, we obtain the expansion coefficients $\breve{d}_{n,i}(E_{b})$ only with $n\leq\mathcal{\color[rgb]{0,0,0}N}_{i}-1$. That is, in our case of $N_{\max}=12$, we have $n\leq 6$ in the ${}^{3}S_{1}$ and $n\leq 5$ in the ${}^{3}D_{1}$ partial waves. Within the HORSE and Efros methods, we obtain coefficients $\breve{d}_{n,i}(E_{b})$ with $n$ ranging up to infinity in each channel. The coefficients $\breve{d}_{n,i}(E_{b})$ with $n>\mathcal{\color[rgb]{0,0,0}N}_{i}-1$ decrease monotonically and are responsible for the description of asymptotic tails of the bound state wave functions $A_{i}e^{-\varkappa r\!}$, where $A_{i}$ is the so-called asymptotic normalization constant in the channel $i$ and $\varkappa=\frac{1}{\hbar}\sqrt{2mE_{b}}$. The description of the asymptotic tails of bound state wave functions is important especially in weakly-bound systems such as the deuteron since they contribute to the rms radius of the system, to its quadrupole moment, to electromagnetic transitions, etc. The asymptotic normalization constants $A_{i}$ are important observables related to the residue of the $S$-matrix pole Baz which are widely used in calculations of various reactions.
V.3 Two-channel resonance
Finally, we consider a resonance in the Noro–Taylor two-channel model originally presented in Ref. Norro1980 which includes two $s$-wave channels ($l_{1}=l_{2}=0$) and the potential
$V_{ij}=\lambda_{ij}r^{2}e^{-r}+E_{i}^{\textrm{thresh.}}\delta_{ij},$ (88a)
where
$\lambda_{11}=1,\quad\lambda_{12}=\lambda_{21}=-7.5,\quad\lambda_{22}=7.5,$ (88b)
and the threshold energies in the channels 1 and 2 are
$E_{1}^{\textrm{thresh.}}=0,\quad E_{2}^{\textrm{thresh.}}=0.1.$ (88c)
The atomic units are used here.
The interaction (88) has a rich spectrum containing four bound states and a number of resonances, and is therefore a good test problem for the numerical validity of a particular approach. We explore the lowest-lying narrow resonance in this system at the energy $E_{r}=4.7682$ au with the width $\Gamma=0.00142$ au.
We present in Fig. 27 the resonant energy $E_{r}$ and width $\Gamma$ obtained by the HORSE method with ${\Omega=1.5}$ au, smoothing parameter $a=5$, and various $N_{\max}$. The energy $E_{r}$ is seen to converge at relatively small $N_{\max}$. In particular, $N_{\textrm{max}}=10$ is sufficient for obtaining $E_{r}$ with an accuracy of 1%. For the width $\Gamma$, approximately $N_{\textrm{max}}=50$ is needed to achieve a similar level of accuracy. Note, however, that the resonance is very narrow and the corresponding $\Gamma$ is small; thus the accuracy of 1% for the width means fixing it within a very small interval of absolute values. For both of these observables, we see a good level of stability beyond $N_{\max}=22$.
The results of calculations of $E_{r}$ and $\Gamma$ within the Efros method with eigenfunction SRFs, $\Omega=1.5$ au, $N_{\max}=20$, and smoothing parameter $a=5$ at various $N$ are presented in Fig. 28 and are compared with the HORSE method result. We see that the level of accuracy within better than 0.1% of the exact result is achieved for $E_{r}$ at $N=10$. For the width $\Gamma$, a relative accuracy of better than 15% is achieved at around $N=19$. We note that both the resonance energy $E_{r}$ and its width $\Gamma$ converge with $N$ to the HORSE results which require larger $N_{\max}$ values to reproduce accurately the width. With $N=19$ we nearly exactly reproduce the HORSE width; thus one needs to perform calculations with larger $N_{\max}$ to describe the exact $\Gamma$ value with higher accuracy.
VI Conclusions
We investigated in detail possibilities of the Efros method VDEfros in describing scattering parameters, wave functions, and in locating $S$-matrix poles to obtain resonant energies and widths or accurate bound state energies and wave functions. To study the convergence and accuracy of the Efros method, we utilized various two-body systems where the exact results may be obtained by other approaches. In particular, we studied the description of $np$ scattering in the ${{}^{1}S}_{0}$ partial wave in the single-channel version of the Efros approach and in the ${{}^{3}SD}_{1}$ coupled partial waves in its two-channel version using the realistic Daejeon16 $NN$ interaction. To investigate the location of $S$-matrix poles, we explored the $\frac{3}{2}^{-}$ resonance in $n\alpha$ scattering described by the Woods–Saxon potential. We located also the $S$-matrix poles in the two-channel systems by calculating the deuteron bound state using the Daejeon16 $NN$ interaction and resonance parameters in the Noro–Taylor two-channel model Norro1980 .
Having in mind future applications in many-body systems in combination with the NCSM NCSMProg , we exploited the harmonic oscillator basis and introduced the Hamiltonian $H^{tr}$ truncated in the oscillator basis at some maximal oscillator quanta $N_{\max}$ as is done in the NCSM. We demonstrated how this truncated Hamiltonian $H^{tr}$ can be utilized within the Efros method. We suggested some modifications of the formalism of the Efros method using achievements developed within the HORSE formalism BangHORSE in scattering theory. The Efros method is equivalent to the HORSE approach if the set of SRFs includes all oscillator functions with oscillator quanta up to $N_{\max}$. This is, however, impractical in many-body applications since such a set of SRFs is extremely large. The great advantage of the Efros method is that it provides a reasonable convergence with significantly reduced SRF sets. One can use as SRFs a limited number of the oscillator functions describing the relative motion in each channel, a number of the eigenfunctions of $H^{tr\!}$, or their mixture. The use of the $H^{tr}$ eigenfunctions seems to be most convenient for future many-body applications of the Efros method in combination with the NCSM. Therefore, most of our studies were performed with the eigenfunction SRF sets. We demonstrated that the smoothing of potential energy matrix elements suggested in Refs. HungOne ; HungTwo significantly improves the convergence of all calculations.
Generally, the Efros method provides accelerated convergence and good accuracy in the description of various observables with a number of eigenfunction SRFs that are accessible in modern NCSM calculation. We did not find problems in calculations of single-channel scattering but in calculations of two-channel scattering, we obtained spikes and discontinuities of some $S$-matrix elements. These shortcomings, however, can be eliminated by reordering and rearranging the SRF set.
We conclude that the Efros method in combination with the NCSM is promising for ab initio studies of elastic scattering and reactions with light nuclei. By locating the $S$-matrix poles within this approach, one can investigate resonant states in light nuclei and derive resonant energies and widths. The $S$-matrix poles associated with bound states which can be calculated within the Efros extension of the NCSM, will improve the results of calculations of bound state energies and wave functions. In particular, it will be possible to describe the asymptotically decreasing tails of the wave functions. These improvements will result in converged and accurate predictions of various nuclear observables like the rms radii, quadrupole moments, asymptotic normalization constants, electromagnetic transition rates, etc. The most important improvements are expected in weakly-bound systems, e. g., in drip-line nuclei which are a focus of modern research in nuclear physics.
Acknowledgments
We are thankful to V. D. Efros for valuable discussions. This work was supported by US DOE Grants DE-SC0023692 and DE-SC0023707 under the Office of Nuclear Physics Quantum Horizons program for the “Nuclei and Hadrons with Quantum computers (NuHaQ)” project.